### v1

# Automatic Fact-Checking with Frame-Semantics

Abstract
We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims, addressing the challenges posed by misinformation in today’s information ecosystem. To support this approach, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data. This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames and data sources from the Organisation for Economic Co-operation and Development (OECD). Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval, indicating a meaningful advancement in automatic fact-checking capabilities. Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future research.
Automatic Fact-Checking with Frame-Semantics
Jacob Devasier, Rishabh Mediratta, Akshith Putta, Chengkai Li University of Texas at Arlington cli@uta.edu
1 Introduction
The proliferation of misinformation presents a significant challenge to the modern information ecosystem. Mitigating the spread of fake news has become a critical concern for researchers, policymakers, and technology developers alike. In response, automatic fact-checking has emerged as an important research area, with the goal of reducing the burden on human fact-checkers by automating key steps in the fact-checking pipeline.
Many previous studies on automatic fact-checking have utilized unstructured data, such as fact-checks from trustworthy sources Wang (2017), to perform claim matching Shaar et al. (2020) and use large language models to provide verdicts and explanations Cheung and Lam (2023b); Singhal et al. (2024a); Khaliq et al. (2024). Some works have also studied structured data, such as tabular data from Wikipedia Chen et al. (2020); Aly et al. (2021) or scientific documents Wang et al. (2021); Akhtar et al. (2022). These works have been studied on both simple claims extracted from Wikipedia Bouziane et al. (2021) and real-world claims Wang et al. (2021); Akhtar et al. (2022).
Previous approaches Ye et al. (2023); Karagiannis et al. (2020); Jo et al. (2019) have found success with using SQL to query Wikipedia tables, which have an average of 13 rows per table Chen et al. (2020). color=green]The last sentence “due to a lack of …” is a bit confusing . I suppose we are talking about Wikipedia tables only. If so, then the notion of “datasets” is moot, because there is only one dataset that’s applicable—all tables in Wikipedia. By “due to a lack of ”, did you intend to say that most tables there are small? — the point I was making is that there are no existing large-scale datasets for structured automatic fact checking, which is why no past works have addressed it Due to a lack of existing datasets, no previous work has studied automatic fact-checking on real-world claims using large-scale structured data. To overcome the limitations of previous methods, we introduce a novel pilot dataset of real-world claims extracted from PolitiFact  [^1] https://www.politifact.com/ annotated for large-scale structured data.
Our proposed pilot dataset serves as the foundation for two case studies focused on leveraging frame semantics for automatic fact-checking. Frame semantics Ruppenhofer et al. (2016); Baker et al. (1998) is a linguistic framework that explores how language encodes meaning through structured representations called frames. These frames capture the essential elements and relationships in various situations, enhancing the understanding of context and intent behind statements.
In the first case study, we utilize the Vote semantic frame Arslan et al. (2020), which represents how an Agent (e.g., a politician) interacts with a particular Issue (e.g., a proposed bill) through voting. Each frame includes specific components known as frame elements, which provide additional context about the roles and relationships involved, such as the voting member, the bill at stake, and the outcome of the vote. color=cyan]This sentence “Each frame includes …” is a little problematic. The previous sentence already discussed Agent and Issue, which are frame elements. And this sentence is a bit redundant with the previous paragraph. Some re-ordering, reorganization may help. This case study includes 79 claims along with their corresponding congressional members and bills, with an average of 4,230 votes per congress member extracted from official U.S. voting records.color=green]you mentioned the frame side of the framework, but not the database side of it. Particularly, the 4,230 votes per member come from the voting record database. Without discussing the database and its role in this, a reviewer may get confused by the mentioning of the number.
The second case study broadens the scope by exploring a variety of additional semantic frames and a diverse collection of datasets from the Organisation for Economic Co-operation and Development (OECD), which contain over 400 tables with an average of 596,000 rows and 13 columns. color=green]should we say “OECD-related claims” instead? Also, “For OECD claims, we annotated 68 claims” is strange. Couldn’t we just say “We annotated 68 OECD-related claims”? We annotated 68 OECD-related claims along with their corresponding frames and mapped each claim to the OECD statistics used to fact-check them. color=green]The notion of “evoke” hasn’t been mentioned. color=green]Similar to the last paragraph, you haven’t discussed the role of the OECD data in fact-checking. These case studies help to identify the challenges associated with automated fact-checking using large-scale structured data.
Our findings demonstrate the value of frame semantics in automatic fact-checking by enabling a structured and explainable understanding of claims. For instance, on voting claims, we found that using frame elements extracted from claims, instead of the entire claim itself, improved evidence retrieval, leading to a +2.1-point increase in recall@10 for voting claims. Similarly, on OECD claims, we saw a +7.3-point increase in recall@5 when frame elements were used to identify relevant OECD tables. These improvements illustrate how frame-semantics color=cyan]earlier you only said “frame semantics”, not parsing. can enhance retrieval by guiding systems toward relevant structured data.
To understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and found a heavy skew towards a few specific frames, including Taking_sides, Speech, Change_position_on_a_scale. These insights guided the selection of frames for our case studies and enable future works to study high-impact frames.
To summarize, our contributions are as follows:
  • We proposed a novel paradigm for automatic fact-checking using frame-semantics.
  • We developed a pilot dataset for automatic fact-checking of real-world claims using large-scale structured data. [^2] https://github.com/anon-naacl25/
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, allowing researchers to target high-impact frames for future studies.
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^3] https://anon-naacl25.github.io/
2 Related Works and Background
Recent advances in automatic fact-checking have been largely driven by the integration of large language models (LLMs) and retrieval-augmented generation (RAG) pipelines. Wang et al. (2024) proposed a unified framework for LLM-based systems which utilizes an internal mechanism to determine which of a selection of LLM-base models should be used to fact-check a particular claim. RAGAR Khaliq et al. (2024) enhances fact-checking by using multimodal inputs and iterative reasoning. Similarly, FactLLaMA Cheung and Lam (2023a) combines pre-trained LLaMA models with external evidence retrieval to verify claims. LLM-Augmenter Peng et al. (2023) integrates external knowledge sources and providing feedback to improve model accuracy. Singhal et al. (2024b) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
For structured data, many approaches have been studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2022) for fact verification. Dater Ye et al. (2023) is a detailed fact verification system which decomposes claims into sub-questions using LLMs to simplify the fact-checking process and is the best-performing model on the TabFact Chen et al. (2020) dataset. In this work, we do not focus on fact verification; however, we choose to utilize LLM-based methods for our fact verification step as they do not require retraining the model when new data is used. color=green]I felt this statement undermines the significance of our work. It focuses on the usage of LLM. But the more important essence of our work is the usage of frames.
2.1 End-to-end Fact-Checking Systems
ClaimBuster Hassan et al. (2017) describes the process of automatic fact-checking as having several core components including a claim matching component, where claims are matched with previous fact checks if they exist, and a knowledge base lookup, which can directly answer questions using internal knowledge, e.g., Wolfram|Alpha. Automatic fact-checking systems, like this work, aim to fulfill the role of the knowledge base. color=cyan]Is that internal or external? More recent studies Guo et al. (2022) have decomposed the process into claim detection, evidence retrieval, and verdict prediction and justification. color=green]Looks strange to have a subsection for a short paragraph. Its connection to the whole section and to the paper’s topic is not clearly explained. — paragraph header looks a bit weirder in my opinion. should we just remove this subsection?
Figure 1: An example of our proposed paradigm. First, the frame and frame elements (FEs) are extracted using a frame-semantic parser (FSP). Then, FEs are mapped to a table which can then be queried to gather evidence for the claim. Finally, the evidence and claim are passed into a fact verification model to check the truthfulness of the claim.
3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment. Figure 1 provides an example of how the system processes the statement “Japan’s life expectancy is the 2nd best in Asia” to illustrate this process.
color=green]Need some better re-organization. the first half of this section has several segments, each with a bold-face title. Then we suddenly start to have subsections for the 2nd half, where each subsection is just a paragraph.
3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing enables the extraction of structured, predefined representations from claims. By identifying frames and their associated frame elements (FEs), the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification. For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.
Frame semantics not only facilitate claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables, other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.
3.2 Data Collection and Evidence Retrieval
The availability of trustworthy data is essential for fact-checking. Rather than relying on search engines that can return inconsistent results or be vulnerable to fake news injections Huynh and Hardouin (2023); Horne et al. (2019), our system stores ground-truth data in an internal database. Using data from reliable sources like the OECD and U.S. Congress ensures the integrity of the evidence.
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy. Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by adapting methods like retrieval-augmented generation Lewis et al. (2020) using text chunking techniques to query large text corpora instead of a database.
3.3 Claim-Evidence Alignment
The final step involves aligning the retrieved evidence with the original claim. Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context. By only using specific spans of the claim, the system can align the claim and evidence more precisely. This approach also enhances explainability, as the system can point to the exact text used in the query.
Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim. For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best. While this claim is not exactly true, the spirit of the claim is true, i.e., that the life expectancy is extremely high in Japan, and thus the model would predict that it is mostly true.color=yellow]This is an intriguing example. But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it? — this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
While fine-tuned approaches have shown slightly better performance on table-based fact-checking Ye et al. (2023), we employ an LLM-based approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions. Because this step is not a key focus of our work, we leave fact verification model-agnostic to allow future improvements. Thus, we can replace the current LLM with any fact verification model which takes structured evidence and a claim to perform fact verification.
4 Fact-Checking Case Studies
In this section we provide two case studies which utilize our proposed paradigm.
4.1 Voting Records
PolitiFact is a leading fact-checking organization that assesses the accuracy of public statements using verifiable evidence. Voting records  [^4] https://www.politifact.com/voting-record/ are a substantial portion of PolitiFact’s fact-checks. In this case study, we focus on automatically fact-checking voting-related claims using the Vote frame and official U.S. congressional records. Specifically, we target the Agent and Issue Frame Elements (FEs), which refer to the voting entity and the topic of the vote, respectively.
Datasets.
We compiled a dataset of official U.S. congressional voting records from the Congress Github Repository. [^5] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.
In their fact-checking process, PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims. In the context of the Vote frame, claims often reference specific congressional bills. To evaluate our system, we constructed an evaluation dataset by extracting all PolitiFact fact-checks made before April 2022 that involve claims related to the Vote frame and reference at least one congressional bill. After manual verification, we collected 79 fact-checks, along with their corresponding bills, to form the evaluation set.
Congress Member Identification.
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use SQL queries to match congress members whose names are similar to the words in the Agent FE. In cases of name ambiguity, e.g., when two members share the same name, we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Although not exhaustive, these lists should cover many common cases. Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz. We handle this by utilizing a list of congress members’ preferred names, supplemented by common alternatives  [^6] https://github.com/carltonnorthern/nicknames when necessary.
Bill Matching.
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing.
Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, which allows us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
Verifying alignment of claim to bill.
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (FE) (for or against). Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”. Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
Claim Verification.
Finally, our system integrates the alignments between each bill and claim to perform fact verification over all of the retrieved evidence. We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix D.4. color=green]Not clear to me what “each individual claim-bill alignement” means.
Figure 2: System design of OECD case study. The frame-semantic parser (orange) facilitates the claim understanding process, table matching and querying (blue) facilitate the evidence retrieval process, and the claim verification step (green) facilitates the fact verification.
4.2 OECD-Related Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics and serves as a strong focal point to explore a wider range of semantic frames. As a visual aid, we include a system overview in Figure 2.
Dataset.
For the OECD Case Study, we collect all of the data tables available on the OECD Data Explorer.  [^7] https://data-explorer.oecd.org/ We construct a SQLite database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totalling 4.1 billion cells.
Relevant Table Identification.
To identify relevant tables for data extraction, we apply semantic similarity models to select the five most similar tables based on the extracted frame elements (FEs). The query consists of one or more FEs derived from frame-semantic parsing, while the target is a specific table. We encode each table’s name and description as text for this purpose. Given the brevity of our queries and the length of the table descriptions, we utilize asymmetric semantic similarity models to ensure precise matches. To minimize the risk of overlooking relevant data, we retrieve the top five most similar tables.color=green]Is “asymmetric semantic similarity” a research field, a technique, or a particular model (maybe in GPT suite or library)? Need to be clear on this and provide link/reference. — a type of model determined by how it is trained, i.e., short query vs long document
Querying Relevant Tables.
Querying the relevant tables presents challenges, as the system must recognize how values in the claim are represented in the database. color=cyan]We said “presents challenges” quite many times in the system. You might want to consider add some variety to the wording. In cases where direct mappings are available (e.g., in the Vote frame case study), we map FEs to specific tables. However, in the OECD case study, where no direct mapping exists between FEs and table columns, the system must predict this mapping. To handle these cases, we employ LLM-based code generation, allowing for the execution of multiple simple queries using programmed logic, rather than relying on complex text-to-SQL generation.color=cyan]Should we include some text-to-SQL references?
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2023), which studied different methods of representing SQL tables for text-to-SQL generation.color=green]What is Gao et al. 2023 about? on the same problem or something else. YOu might want to make this clear. Alongside the schema, we include representative example values from the database for each column. color=green]“Columns …. include all values” doesn’t seem correct. — referring to how they are encoded into text Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values. For columns with a moderate number of unique values, we use a semantic similarity model to select the top ten most relevant values based on the claim. Given the large number of cells generated by this process, we use an LLM to refine the query and filter out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompt for this LLM can be found in Appendix D.3. color=green]Is this semantic similarity model the same as the aymmetric one mentioned in the last section? If yes, make it clear; if not, need reference/link, based on the same concern mentioned earlier. — each model is different because they perform differently on the data, the selection is discussed in the experiments color=green]Are we missing some details here about how we used LLM? Are there prompts?
Fact Verification
In the final step, we used an LLM-based fact verification model. We represent the extracted evidence as a list of tab-separated values based on the data retrieved in the previous step. The model is instructed using the prompt in Appendix D.4 and outputs one of five verdicts: false, mostly false, half-true, mostly true, or true.color=cyan]If time allows, explain the connection to PolitiFact’s Truth-o-meter.
5 Experiments
5.1 Datasets
Fact-checking Frames.
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in this dataset along with the newly-defined frames. These frames can be found in Appendix 6.
PolitiFact Fact-checks
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles collected as of April 2022. Each article contains a claim, a detailed fact-check, a verdict, and a list of sources. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we identified 79 fact-checks that cite congress.gov and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset.
color=cyan]PolitiFact fact-checks and PolitiFact fact-check articles. Mention very close to each other. color=green]the claim cites congress.gov? color=green]“bill cited was referenced by the claim”. cited where? also the claim would reference a bill?
Additionally, we collected 68 articles that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the statistics which can be used to fact-check it. These two sets of claim-evidence pairs—the voting-related claims and the OECD-related claims—serve as evaluation sets for our case studies. color=green]these articles cite the website? color=green]“To collect … we collect” color=green]Which two sets?
Model Frames Frame Acc FE Acc
Random Vote 0.488 0.254
GPT-4o-mini Vote 0.974 0.618
Vote FSP Vote 0.990 0.889
Random OECD 0.602 0.000
GPT-4o-mini OECD 0.537 0.372
GPT-4o-mini* OECD 0.713 0.461
OECD FSP OECD 0.742 0.873
Table 1: Performance of frame-semantic parsing model on fact-checking frames alongside simple baseline.
Frame Samples (%)
Taking_sides 7,152 (34.0%)
Speech 6,010 (28.6%)
Change_position_on_a_scale 5,547 (26.4%)
Comparing_two_entities 5,530 (26.3%)
Cause_change_of_position_on_a_scale 4,675 (22.2%)
Vote 3,229 (15.4%)
Comparing_at_two_different_points_in_time 2,436 (11.6%)
Conditional_occurrence 2,355 (11.2%)
Creating 2,194 (10.4%)
Occupy_rank 1,106 (5.3%)
Oppose_and_support_consistency 1,010 (4.8%)
Recurrent_action_in_Frequency 935 (4.4%)
Ratio 932 (4.4%)
Capability 869 (4.1%)
Occupy_rank_via_superlatives 767 (3.6%)
Uniqueness_of_trait 497 (2.4%)
Occupy_rank_via_ordinal_numbers 329 (1.6%)
Recurring_action 187 (0.9%)
None 12 (0.1%)
Table 2: Distribution of semantic frames in PolitiFact.
5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combine the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unify frame and frame element identification into a single model. This model is then trained only on the fact checking frames introduced in Arslan et al. (2020).
To evaluate the performance of our model for the case studies, we compare it with a GPT-4o-mini model using OpenAI’s structured generation [^8] At the time of submission, GPT-4o-mini is the only model which supports structured generation.. We evaluate each model using accuracy on frame identification and exact match accuracy on FE identification in Table 1. We define a separate prompt for each case study in Appendix D.1.
5.3 Claim Coverage
One concern which may arise is regarding the applicability of our system for assisting in the fact checking process. To understand this, we aim to identify the coverage of a frame-based approach towards automatically fact-checking claims. Due to the lack of annotations for all of the fact-checking frames in Arslan et al. (2020), we utilize a zero-shot GPT-4o-mini model to identify the frames evoked in the claim of each PolitiFact fact-check (Table 2). Claims which do not evoke a frame studied in this work are categorize as none. The prompt used for this model can be found in Appendix D.1.
Model Data R@K
BM25 OECD 0.323
distilbert-tas-b OECD 0.505
text-embedding-3-large OECD 0.610
roberta-base-v2 OECD 0.726
text-embedding-3-large Vote 0.032
stella_en_1.5B_v5 Vote 0.144
distilbert-tas-b Vote 0.165
Table 3: Performance comparison of different semantic similarity models on extracted frame elements. K=10 for Vote and K=5 for OECD.
Model Query Data R@K
RoBERTa (v2) Full claim OECD 0.653
RoBERTa (v2) FE OECD 0.726
Max Possible - OECD 0.910
distilbert-tas-b Full claim Vote 0.143
distilbert-tas-b Issue FE Vote 0.165
Max Possible - Vote 0.568
Table 4: Performance of different methods of representing the query in the OECD case study’s table matching. OECD and Vote claims use Recall@5/10, respectively.
5.4 Evidence Retrieval
We study the capability of our system to retrieve the relevant information which can be used to fact-check the claim. To do this, we found the percentage of bills and statistics available in our database for voting and OECD-related claims, respectively. This serves as a performance ceiling on our collected claims as our system would be unable to fact-check claims which do not have available data. These values are found in Table 4 (Max Possible).
For semantic similarity, we compare commonly used embedding models (Table 3), namely RoBERTa Liu et al. (2019) and DistilBERT-TAS-B Hofstätter et al. (2021). Additionally, we compare with OpenAI’s text-embedding-3-large model and provide a baseline BM25 approach. We also experiment with different approaches of representing the query used to find relevant data. In Table 4, we compare the performance of using the entire claim with only the relevant FE. To evaluate the performance of these queries, we calculate the recall@K, where K=5 for OECD claims and K=10 for Vote claims. The relevant FEs used for each frame can be found in Appendix A.
Model Dataset Accuracy
GPT-4o Naive Vote 0.044
Our w/ Irrelevant Vote 0.076
Our w/o Irrelevant Vote 0.207
Our w/ Irrelevant ± Vote 0.253
GPT-4o Naive ± Vote 0.324
Our w/o Irrelevant ± Vote 0.690
GPT-4o Naive OECD 0.073
Our w/ Irrelevant OECD 0.214
Our w/o Irrelevant OECD 0.429
GPT-4o Naive ± OECD 0.537
Our w/ Irrelevant ± OECD 0.607
Our w/o Irrelevant ± OECD 0.821
Table 5: Fact verification performance on different case studies. ± indicates the off-by-one performance. Bolded values indicate the best-performing settings overall and underlined values are only considering exact matches.
5.5 Fact Verification
Finally, we study the ability of our fact verification system to fact-check claims using structured data by comparing the verdicts presented in each PolitiFact fact-check to the model’s predicted verdict. We have replaced PolitiFact’s “pants on fire” verdicts with a false verdict. To determine whether the LLM is relying on internal knowledge or if it is using the provided data, we evaluate the performance of GPT-4o without any data. We measure accuracy as the primary evaluation metric and also calculate off-by-one accuracy, where, for example, a prediction of “mostly false” is considered correct if the label is “false.” The results of these experiments are summarized in Table 5.
6 Results
6.1 Frame-Semantic Parsing
We evaluated the frame-semantic parsing model on fact-checking frames introduced by Arslan et al. (2020), comparing it against GPT-4o-mini due to its structured output capability. Table 1 shows the performance on both the Vote and OECD frames. We found that the GPT-4o-mini model tends to over-predict the number of frames in a sentence. When we only use the first predicted frame, the performance is increased significantly. We saw similar cases with frame element predictions where GPT-4o-mini tends to predict more frame elements than actually exist in the input. We include examples of each of these cases in Appendix C.2.
6.2 PolitiFact Survey
Similar to Section 6.1, we found that GPT-4o-mini tends to predict more frames than there may actually be. We believe this is likely due to the annotations from Arslan et al. (2020) being sampled from PolitiFact and likely having a similar structure. We found the annotated samples evoked 1.1 frames on average while GPT-4o-mini predicts an average of 2.1 frames. Another possible explanation is that the model is able to identify frames for lexical units– words which evoke frames–which are not defined.
One indication of this is on the Vote frame, which only has a lexical unit defined for vote.v; however, it is possible to evoke the Vote frame using other words. For example, the statement “I passed a bill” implies that a vote of affirmation was cast on that bill. We found only 1,198 (5.7%) PolitiFact fact-checks which mention “vote” while the model predicts that almost 3 times as many evoke the Vote frame. While we found many claims which evoke a Vote frame without an explicit mention of the word “vote”, further analysis is needed to get a more accurate understanding of the true distribution of frames evoked by factual claims.
6.3 Evidence Retrieval
We found that embedding models’ performance is task-specific and there is no single-best embedding model for every task. For example, RoBERTa performed the best for finding the right OECD table while DistilBERT-TAS-B was better for matching claims to their corresponding bills. We also found that, in general, using the claim-specific frame element performs better than using the entire claim as a query, as shown in Table 4.
For voting-related claims, we found that identifying the bill referenced in a particular claim is very difficult. Our best performing model, DistilBERT-TAS-B, was only able to obtain a Recall@10 of 0.165. For OECD-related claims, identifying the table to use to fact-check the given claim was much easier, with the best model, RoBERTa, achieving a recall@5 of 0.726. This is likely because the similarity between the relevant FE and the table names were better-represented by semantic similarity models. This indicates a potential area of significant improvement in the methodology and representation of bill texts according to their topics and implications, in addition to the text itself.
6.4 Fact Verification
On OECD claims, 62% of the generated queries were able to retrieve relevant data while only 36% of the voting claims were able to retrieve relevant bills. These values put an upper bound on the capabilities of our fact verification model as without evidence, the model will not be able to provide useful fact-checks. In general, we found that language models tend not to predict a claim to be true based on the given data, yet they have no issue with predicting a claim to be false, and often do so. We believe this is analogous to proofs, it is quite easy to prove by contradiction, if you find a single vote or country statistic which contradicts the claim, while it is much harder to assert that a claim is true without looking at all of the data.
We found that on Vote claims, a naive GPT-4o correctly predicted the truthfulness of the factual claim with an accuracy of 0.286; however, when voting data is added, the performance drops to 0.076. This suggests that, while there may be some internal knowledge on certain people and bills, it is unlikely that the model is relying on it when given specific voting records. This also indicates that the prompt given to the model is sufficient in preventing the LLM from relying on internal knowledge.
When excluding claims for which data was not found, we observed absolute performance increases of +13.1% and +21.5% for Vote and OECD claims, respectively. We also found that, while the exact-match accuracy of the system is low, there is a strong correlation between predictions and labels, as indicated by the off-by-one accuracy in Table 5. We include example outputs for the explanations of these models in Appendix C.1.
7 Conclusion
This research contributes to the evolving field of automatic fact-checking by proposing a novel paradigm that incorporates frame semantics to enhance the evidence retrieval process and enable fine-grained control of downstream tasks. By developing and presenting a pilot dataset specifically annotated for real-world claims, this work helps to address the underexplored challenge of fact-checking using massive, structured data sources.
Our findings demonstrate the promise of frame-semantic parsing in improving evidence retrieval for automatic fact-checking. In both case studies, we observed that using frame elements extracted from claims instead of entire claims led to substantial improvements in recall metrics, highlighting the utility of frame-based understanding in identifying relevant evidence. These results suggest that semantic frames not only offer a structured and explainable approach to parsing claims but also guide retrieval models more effectively toward the correct information within large datasets.
Looking ahead, the integration of frame semantics into fact-checking systems holds promise for further advancements in both accuracy and interpretability. Future work could explore broader applications of this approach, including expanding the dataset, refining the models, and incorporating additional semantic frames to cover more claims.
Limitations
Despite the advancements made in this study, our system’s limitations should be acknowledged. While this study focuses primarily on improving evidence retrieval and frame-semantic parsing, it does not comprehensively evaluate different fact verification approaches. Fact verification is critical for automatic fact-checking, and a deeper investigation into various verification strategies, including fine-tuned models and LLM-based approaches, could further enhance system performance.
Another limitation is that our system’s effectiveness is constrained by the availability of reliable, manually curated data. For claims that lack relevant data in our databases (e.g., OECD tables or U.S. congressional records), the system cannot retrieve evidence, which limits its fact-checking capabilities. This limitation demonstrates the importance of expanding data sources and improving coverage in future iterations of the system.
Ethics and Risks
Automated fact-checking systems, such as the one presented in this work, bring both opportunities and ethical challenges. One key concern is the potential spread of misinformation due to model limitations or errors. Users may over-rely on AI verdicts, leading to the amplification of false positives or negatives, especially in politically sensitive contexts. Biases present in both the models (e.g., GPT-4o-mini, RoBERTa) and the datasets (e.g., PolitiFact, OECD data) could skew fact-checking outcomes, favoring certain political narratives or ideologies.
Another concern is the transparency and accountability of AI systems. If users are unaware of how models arrive at their conclusions, it may be difficult to hold these systems accountable for erroneous outcomes. This opacity could diminish trust in both the fact-checking tool and the institutions that deploy it. Moreover, the collection and processing of political data raise potential privacy concerns, especially regarding the use of public records in ways individuals may not expect.
Marginalized communities may also be disproportionately affected by such systems, as they might misinterpret claims relevant to those groups or lack sufficient representation in training data. Similarly, the system could be exploited by adversaries who craft ambiguous or misleading claims designed to confuse AI, leading to manipulated fact-checks.
Lastly, the over-dependence on specific sources like PolitiFact or congress.gov could limit the tool’s scope. Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
Appendix A Studied Frames
Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings.
Frame Frame Element Frame Definition
Occupy_rank Dimension Items occupying a certain Rank within a hierarchy.
Occupy_rank_via_superlatives Dimension An Item occupying a Rank specified by a superlative.
Comparing_two_entities Comparison_criterion Comparing two entities using a Comparison_criterion and qualifying with a Degree.
Comparing_at_two_different_ points_in_time Comparison_criterion Comparing an Entity with itself at two different points in time using a Comparison_criterion and qualifying with a Degree.
Cause_change_of_position_ on_a_scale Item Words that indicate an Agent or Cause affects the position of an Item on a scale.
Capability Event An Entity meets the pre-conditions for participating in an Event.
Table 6: Frames used for OECD Case Study, along with the Frame Elements extracted from corresponding Frames.
Appendix B Reproducibility
We trained our frame-semantic parser using a NVIDIA GTX 1080 Ti in about 6 hours on a single run without any hyperparameter tuning. Github Copilot was used in the development of some of our code.
B.1 Software License
Our source code and data are released with the MIT License.
Model Parameters URL
msmarco-distilbert-base-tas-b 66M msmarco-distilbert-base-tas-b
msmarco-roberta-base-v2 125M msmarco-roberta-base-v2
msmarco-roberta-base-v3 125M msmarco-roberta-base-v3
stella_en_1.5B_v5 1.5B stella_en_1.5B_v5
text-embedding-3-large unknown text-embedding-3-large
Table 7: Models and Their Corresponding Hugging Face URLs
B.2 Models
Unless stated otherwise, all models used are the “base” variant. All semantic similarity models are available via the sentence-transformers library. [^9] https://huggingface.co/sentence-transformers Table 7 contains all of the models used and their details.
Claim Predicted Verdict Explanation Fact Check
Americans work way more than an average of industrialized countries around the world. Mostly True The data shows that the United States has an annual average of 1829 hours worked per person, which is higher than many other industrialized countries in the dataset, such as Germany (1400.4 hours), France (1518.0 hours), and Japan (1729.0 hours). However, there are some countries like Mexico (2229.0 hours), Korea (2076.0 hours), and Chile (1994.0 hours) with higher working hours than the U.S. This indicates that while Americans work more than many other industrialized countries, they do not work more than all. The claim is mostly true but lacks the nuance that there are industrialized countries where people work more hours. PolitiFact
The United States is the ONLY industrialized country without universal healthcare. True The retrieved data indicates that countries such as Italy, the Netherlands, Spain, Canada, Korea, Australia, France, Germany, the United Kingdom, and Sweden all have near-universal health insurance coverage through government or compulsory schemes, with coverage rates often very close to 100% of the population. In contrast, the United States has only 38.1% of its population covered by government or compulsory health insurance, suggesting that it lacks universal healthcare coverage. This supports the claim that the United States is the only industrialized nation without such a system based on the data provided. PolitiFact
Table 8: OECD fact check example predictions and explanations
Appendix C Examples
In this section, we provide a few example outputs of our system to give a better understanding of the outputs of our models.
Claim Bill Title Vote Alignment Prediction
Marco Rubio voted against the bipartisan Violence Against Women Act 112 S 1925 Violence Against Women Reauthorization Act of 2012 Nay Supports True
17 HR 3233 National Commission to Investigate the January 6 Attack on the United States Capitol Complex Act Nay Irrelevant
117 HR 350 Domestic Terrorism Prevention Act of 2022 Nay Irrelevant
Chuck Grassley was voting to slash Medicare when voting against the debt ceiling bill 117 S 610 Protecting Medicare and American Farmers from Sequester Cuts Act Nay Supports Mostly False
117 S 1301 Promoting Physical Activity for Americans Act Nay Irrelevant
17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
Table 9: Examples of voting-related claims with the corresponding retrieved bills, votes on the bill, the vote-claim alignment, and fact verification prediction.
C.1 Vote-Claim Alignment
We include two sample claims and their corresponding top three retrieved bills along with the vote-claim alignment for each bill in Table 9.
C.2 Fact Verification Predictions
Vote
We include two claims and their corresponding predictions based on the retrieved bills in Table 9.
OECD
We include two predictions and their explanations on OECD claims in Table 8.
Appendix D LLM Prompts
D.1 Frame-Semantic Parsing
Listing 1: Frame-semantic parser for the Vote frame.
Identify if a Vote semantic frame is evoked in a given sentence. If it is, extract and list the relevant frame elements associated with the voting event.
A Vote frame is defined as: An Agent makes a voting decision on an Issue.
Frame elements to identify:
- Agent: The conscious entity (usually a person) executing the voting decision.
- Issue: The subject or matter that the Agent is voting on with a particular position.
- Side: Additional person(s) involved in the voting on the same Issue alongside the Agent.
- Position: The stance the Agent takes, expressing whether their vote is in favor or against the Issue.
- Frequency: How often the Agent has made this voting decision regarding the Issue.
- Time: When the Agent performed the voting decision.
- Place: Location where the vote took place.
- Support_rate: The percentage of the Agents total votes aligning with the Side over a set period.
# Notes
- If a frame element is not mentioned in the sentence, use "" for its value.
- Carefully distinguish between elements; some may have overlapping characteristics.
- Consider synonyms or variations of terms related to voting when analyzing the sentence.
- Frame elements should quote exactly from the input
Listing 2: Frame-semantic parser for OECD frames.
Identify semantic frames evoked by a given factual claim and extract relevant frame elements for each identified frame.
Consider the predefined semantic frames and their elements:
- Occupy_rank_via_superlatives:
- Item: Entity occupying the rank.
- Rank: Rank held, often defined by a superlative.
- Dimension: Aspect along which the ranking occurs (e.g., speed, age).
- Comparison_set: Group of entities being compared.
- Time: Time period when the item occupies the rank.
- Occupy_rank:
- Item: Entity occupying the rank.
- Rank: Rank held.
- Dimension: Criterion used for ranking.
- Comparison_set: Set of entities being ranked.
- Time: Time period when the item holds the rank.
- Change_position_on_a_scale:
- Item: Entity whose position on a scale changes.
- Attribute: Property or scale of the change.
- Difference: Amount of change.
- Final_value: Final position on the scale.
- Initial_value: Initial position on the scale.
- Path: Progression between positions on the scale.
- Speed: Rate of change.
- Correlated_variable: Related variable that changes with the Attribute.
- Manner: Method of performing the action.
- Degree: Extent to which the change occurs.
- Circumstances: Context or conditions for the change.
- Result: Outcome of the change.
- Group: Collection of Items undergoing change.
- Time: Time frame during which the change occurs.
- Duration: Length of time for the change to happen.
- Value_range: Range of values along the scale.
- Initial_correlate: State corresponding to the Initial_value.
- Final_correlate: State corresponding to the Final_value.
- Place: Location where the Attribute is measured.
- Initial_state: State of the Item before the change.
- Final_state: State of the Item after the change.
- Period_of_iterations: Duration from start to stop of repeated changes.
- Particular_iteration: Specific instance of a repeated event.
- Containing_event: Broader event encompassing the change.
- Explanation: Rationale for the change.
- Comparing_two_entities:
- Entity_1: First entity in comparison.
- Entity_2: Entity serving as the comparison point.
- Comparison_criterion: Criterion for comparison.
- Degree: Extent of the comparison.
- Time: Period over which the comparison takes place.
- Comparing_at_two_different_ points_in_time:
- Entity: Entity compared to itself at different times.
- First_point_in_time: First time period in comparison.
- Second_point_in_time: Second time period in comparison.
- Comparison_criterion: Criterion being compared.
- Degree: Extent of the difference.
- Difference: Magnitude of change between the two times.
# Steps
1. Carefully read and analyze the factual claim.
2. Determine which semantic frame(s) are invoked by the claim.
3. For each evoked frame, extract the appropriate frame elements directly from the claim.
4. Document each evoked frame and the extracted elements.
# Output Format
- List each evoked frame followed by its extracted elements in a structured form.
- Ensure all extracted frame elements are clearly labeled and match the exact inputs.
# Notes
- Some claims may trigger multiple frames; ensure all applicable frames are evaluated.
- If specific frame elements cannot be determined, note their absence for completeness.
- If a frame element does not explicitly appear in the claim, leave it blank.
D.2 Claim-Bill Alignment
Listing 3: Claim-Bill Alignment
Given the following factual claim, bill summary, and vote on the bill, evaluate whether the content of the bill summary and the voting record align with the given claim. You may consider factors such as the main objectives of the bill and unintended or implicit consequences. Your task is to determine if the information provided in the bill summary and the voting record supports or refutes the given factual claim. Return your explanation and one of the following labels in JSON format.
Bill Summary:
{summary}
Vote: {vote_type}
Claim: {claim}
Labels:
Supports - The vote on this bill directly or indirectly supports the claim.
Refutes - The vote on this bill explicitly refutes the claim.
Inconclusive - The vote on this bill does not provide enough information to support or refute the claim.
Irrelevant - The vote on this bill is not relevant to the claim at all.
D.3 OECD Data Query
Listing 4: Retrieve Data for Fact-Checking Claim
Your task is to write a Python function named retrieve_data() that retrieves data for fact-checking the claim:
{claim}
Given the following database schemas from OECD_Data.db:
{table_descriptions}
Instructions:
- Use LIKE only for textual columns. For numerical columns (e.g., year, value), use appropriate comparison operators like = or <=.
- Do not modify the database.
- The function must contain all necessary imports inside it and take no parameters, though passing parameters (e.g., file paths or claim details) should be considered in future iterations.
- Return the data in pandas DataFrames. Multiple DataFrames can be returned as a list if needed.
- If no relevant data is found, return Data is not Available’.
- Use only actual values found in the columns. If aggregation is possible (e.g., summing or averaging categories), return aggregated results, unless the claim specifies otherwise. Use appropriate aggregation methods based on the data.
- Always return relevant columns (avoid ’*’ when possible). Select the necessary columns based on the claim.
- Treat we in the claim as referring to United States’, but consider the context of the claim for potential exceptions.
- If any unit_of_measure is National currency’, use the USD_value column instead of value’.
- Use the nearest available date based on the provided dictionary: {nearest_dates}. If no close date is found, return Data is not Available’.
- For multiple tables, create multiple queries as needed. If combining data from multiple tables is required, ensure consistency in merging and handling differences in structure.
- Do not filter by country or unit_of_measure’; that will be handled later.
Output:
- Return a list of pandas DataFrames or Data is not Available if no relevant data is found.
Listing 5: Clean Data for Fact-Checking Claim
Your task is to write a Python code function named clean_data()‘ to filter and clean data for fact-checking the claim {claim} from OECD_Data.db‘.
Extracted dataframes from the previous step:
{list_of_dfs}
These dataframes were generated using this code:
{code}
The schema for these dataframes is:
{schema_str}
Notes:
- The function must have all imports inside it and take no parameters.
- This function will run independently of the given code, meaning you need to re-extract the data from the database to clean it for fact-checking the claim:
{claim}
- Do not modify the original way the data was extracted unless the input code is very incorrect; just add more filters/conditions based on the claim.
- Keep the same aggregated structure from the original extraction code unless the claim specifies otherwise.
- Filter by the nearest date: {nearest_date}.
- Exclude tables only if they are irrelevant to the claim.
- Use only the actual tables and data that were extracted; do not make up new tables.
- Add more filters to narrow down to claim-relevant metrics; try to have one value per non-numeric column wherever possible.
- Claim-relevant metrics may appear in columns other than ‘’measure’‘, so reference the claim to identify the appropriate columns to filter by.
- Use the retrieval code to identify useful information for your filters, if needed.
- When filtering by ‘’unit_of_measure’‘, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
Output:
- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return ‘’Data is not Available’‘.
D.4 Fact Verification
Listing 6: Vote - Verify Claim and Provide Verdict
You are given the following claim, and 5 bills along with the bill title, a short summary of the bill, the vote cast on the bill and the alignment of the bill with the claim.
Alignment here means whether an individual bill supports, refines, or is irrelevant to the claim.
Your task is to give one of the given fact-check labels to the claim based on the evidence which are the bills.
You may consider factors such as the main objectives of the bill and unintended or implicit consequences.
Return your explanation and one of the following labels in JSON format.
Claim: {claim}
Bill Title 1: {bill_title_1}
Bill Summary 1: {bill_summary_1}
Vote 1: {vote_type_1}
Alignment1: {alignment_1}
Bill Title N: {bill_title_N}
Bill Summary N: {bill_summary_N}
Vote N: {vote_type_N}
Alignment N: {alignment_N}
Labels:
True: The given bills support the claim.
MostlyTrue: The given bills mostly support the claim.
HalfTrue: The given bills can only partly support or refute claim.
MostlyFalse: The given bills mostly refute the claim.
False: The given bills refute the claim.
Irrelevant: The given bills are not relevant to the claim.
Return the JSON object with the label and the explanation. The fields should be Label and Explanation’.
Listing 7: OECD - Verify Claim and Provide Verdict
Your task is to verify the claim using the retrieved data and provide a verdict.
Claim: {claim}
Retrieved Data:
{formatted_data}
Instructions:
- Use only the provided data, regardless of its perceived relevance, to analyze the claim.
- The verdict must be based solely on the retrieved data. Do not rely on external knowledge.
- Ensure you consider the spirit of the claim in the fact-check and not just the precise verbage.
Verdict Categories:
- True: The statement is fully accurate, with no significant information missing.
- Mostly True: The statement is accurate but requires clarification or additional context.
- Half-True: The statement is partially accurate but omits important details or misrepresents the context.
- Mostly False: The statement contains some truth but overlooks key facts that would significantly alter the impression.
- False: The statement is completely inaccurate.
Output:
- Provide a verdict in the format:
Verdict: [False, Mostly False, Half-True, Mostly True, True]; Explanation: [Your reasoning].

### v2

@@
-Automatic Fact-Checking with Frame-Semantics
+Task-Oriented Automatic Fact-Checking with Frame-Semantics
 Abstract
-We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims, addressing the challenges posed by misinformation in today’s information ecosystem.
-To support this approach, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data.
-This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames and data sources from the Organisation for Economic Co-operation and Development (OECD).
-Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval, indicating a meaningful advancement in automatic fact-checking capabilities.
-Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future research.
-Automatic Fact-Checking with Frame-Semantics
+We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims and guide the process of fact-checking them.
+To support this, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data.
+This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames based on data sources from the Organisation for Economic Co-operation and Development (OECD).
+Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval and explainability for fact-checking.
+Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future work in this direction.
+Task-Oriented Automatic Fact-Checking with Frame-Semantics
 Jacob Devasier, Rishabh Mediratta, Akshith Putta, Chengkai Li University of Texas at Arlington cli@uta.edu
 1 Introduction
-The proliferation of misinformation presents a significant challenge to the modern information ecosystem.
-Mitigating the spread of fake news has become a critical concern for researchers, policymakers, and technology developers alike.
-In response, automatic fact-checking has emerged as an important research area, with the goal of reducing the burden on human fact-checkers by automating key steps in the fact-checking pipeline.
-Many previous studies on automatic fact-checking have utilized unstructured data, such as fact-checks from trustworthy sources Wang (2017), to perform claim matching Shaar et al.
-(2020) and use large language models to provide verdicts and explanations Cheung and Lam (2023b); Singhal et al.
-(2024a); Khaliq et al.
+The proliferation of misinformation poses a critical challenge to the modern information ecosystem, threatening informed decision-making and public trust.
+Addressing this challenge requires scalable and reliable methods to verify the accuracy of claims—particularly novel ones that human fact-checkers may not have previously encountered.
+Automatic fact-checking has thus emerged as a crucial research area, aiming to reduce the manual burden of verifying claims by automating key steps in the fact-checking pipeline.
+Much of the previous work on automatic fact-checking has focused on unstructured data, such as fact-checks from trustworthy sources, to perform claim matching Shaar et al.
+(2020) and use large language models to produce fact-check verdicts and explanations Cheung and Lam (2023b); Singal et al.
+(2024); Khaliq et al.
 (2024).
-Some works have also studied structured data, such as tabular data from Wikipedia Chen et al.
+Structured tabular data from Wikipedia Chen et al.
 (2020); Aly et al.
-(2021) or scientific documents Wang et al.
+(2021) and scientific documents Wang et al.
 (2021); Akhtar et al.
-(2022).
-These works have been studied on both simple claims extracted from Wikipedia Bouziane et al.
+(2022) have also been utilized to vet synthetic claims extracted from Wikipedia Bouziane et al.
 (2021) and real-world claims Wang et al.
@@
 (2022).
-Previous approaches Ye et al.
-(2023); Karagiannis et al.
-(2020); Jo et al.
-(2019) have found success with using SQL to query Wikipedia tables, which have an average of 13 rows per table Chen et al.
-(2020).
-color=green]The last sentence “due to a lack of …” is a bit confusing .
-I suppose we are talking about Wikipedia tables only.
-If so, then the notion of “datasets” is moot, because there is only one dataset that’s applicable—all tables in Wikipedia.
-By “due to a lack of ”, did you intend to say that most tables there are small?
-— the point I was making is that there are no existing large-scale datasets for structured automatic fact checking, which is why no past works have addressed it Due to a lack of existing datasets, no previous work has studied automatic fact-checking on real-world claims using large-scale structured data.
-To overcome the limitations of previous methods, we introduce a novel pilot dataset of real-world claims extracted from PolitiFact  [^1] https://www.politifact.com/ annotated for large-scale structured data.
-Our proposed pilot dataset serves as the foundation for two case studies focused on leveraging frame semantics for automatic fact-checking.
-Frame semantics Ruppenhofer et al.
-(2016); Baker et al.
-(1998) is a linguistic framework that explores how language encodes meaning through structured representations called frames.
-These frames capture the essential elements and relationships in various situations, enhancing the understanding of context and intent behind statements.
-In the first case study, we utilize the Vote semantic frame Arslan et al.
-(2020), which represents how an Agent (e.g., a politician) interacts with a particular Issue (e.g., a proposed bill) through voting.
-Each frame includes specific components known as frame elements, which provide additional context about the roles and relationships involved, such as the voting member, the bill at stake, and the outcome of the vote.
-color=cyan]This sentence “Each frame includes …” is a little problematic.
-The previous sentence already discussed Agent and Issue, which are frame elements.
-And this sentence is a bit redundant with the previous paragraph.
-Some re-ordering, reorganization may help.
-This case study includes 79 claims along with their corresponding congressional members and bills, with an average of 4,230 votes per congress member extracted from official U.S.
-voting records.color=green]you mentioned the frame side of the framework, but not the database side of it.
-Particularly, the 4,230 votes per member come from the voting record database.
-Without discussing the database and its role in this, a reviewer may get confused by the mentioning of the number.
-The second case study broadens the scope by exploring a variety of additional semantic frames and a diverse collection of datasets from the Organisation for Economic Co-operation and Development (OECD), which contain over 400 tables with an average of 596,000 rows and 13 columns.
-color=green]should we say “OECD-related claims” instead?
-Also, “For OECD claims, we annotated 68 claims” is strange.
-Couldn’t we just say “We annotated 68 OECD-related claims”?
-We annotated 68 OECD-related claims along with their corresponding frames and mapped each claim to the OECD statistics used to fact-check them.
-color=green]The notion of “evoke” hasn’t been mentioned.
-color=green]Similar to the last paragraph, you haven’t discussed the role of the OECD data in fact-checking.
-These case studies help to identify the challenges associated with automated fact-checking using large-scale structured data.
-Our findings demonstrate the value of frame semantics in automatic fact-checking by enabling a structured and explainable understanding of claims.
-For instance, on voting claims, we found that using frame elements extracted from claims, instead of the entire claim itself, improved evidence retrieval, leading to a +2.1-point increase in recall@10 for voting claims.
-Similarly, on OECD claims, we saw a +7.3-point increase in recall@5 when frame elements were used to identify relevant OECD tables.
-These improvements illustrate how frame-semantics color=cyan]earlier you only said “frame semantics”, not parsing.
-can enhance retrieval by guiding systems toward relevant structured data.
-To understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and found a heavy skew towards a few specific frames, including Taking_sides, Speech, Change_position_on_a_scale.
-These insights guided the selection of frames for our case studies and enable future works to study high-impact frames.
+However, these studies only consider data which have already been processed and prepared for easy consumption by readers.
+This reliance on highly-curated data hinders their ability to fact-check novel claims which do not have such clean evidence readily available.
+To bridge this gap our work proposes a novel pilot dataset of real-world factual claims related to two high-volume structured databases on U.S.
+congressional voting records and country statistics.color=yellow]Saying we created the dataset to fill the gap gives me the feeling of we invent the nail.
+Instead, consider saying the verification of many claims rely on data that only exist in much larger tables.
+These databases significantly differ from past works in both volume and detail of their tables.
+For example, Wikipedia tables consist of an average of 13 rows Chen et al.
+(2020) while our database of country statistics has an average of 596,000 rows across over 400 tables.
+This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed, even for those solutions designed for large tables Nahid and Rafiei (2024).
+Fact-checking novel claims requires fine-grained methodology to understand the claim and to map important parts of the claim to their relevant data.
+Furthermore, transparency and explainability are critical in fact-checking.
+We identify frame semantics Baker et al.
+(1998) as a viable solution which addresses each of these points.
+Frame semantics is a linguistic framework that explores how meaning is encoded through structured representations called frames.
+These frames capture the important elements and relationships in various situations and enable explainable understanding of claims and extraction of their key parts.
+The existence of specific frames in a claim also enables fine-grained task-specific handling of claims for fact-checking.
+Our proposed pilot dataset serves as the foundation for two case studies that demonstrate how frame semantics can be used to guide specific fact-checking processes and enhance the explainability of automatic fact-checking.
+In the first case study, we focus on the Vote frame Arslan et al.
+(2020), which models how an Agent (e.g., a legislator) interacts with an Issue (e.g., a bill) through the act of voting and optionally takes a Position (for or against) on the Issue.
+These Issue and Agent elements are then used to implement voting-specific fact-checking methods.
+The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
+To evaluate these case studies, we collected 79 and 68 real-world claims for the voting records and OECD statistics, respectively.
+Each claim is manually annotated for its particular frames and fact-check verdict retrieved from PolitiFact.
+Our case studies demonstrate that frame semantics offers a valuable framework for task-specific fact-checking.
+Frame-semantics can also offer additional performance benefits for particular tasks.
+For example, on Vote and OECD-related claims, retrieving relevant evidence by querying with extracted task-specific frame elements outperformed using the entire claim by 14% and 11%, respectively.
+The performance of particular retrieval methodologies also significantly varied between the two case studies, indicating a need to select different approaches for each specific task to achieve optimal performance.
+Furthermore, instructing large language models (LLMs) to fact-check claims benefits from different prompts depending on the type of claim.
+These factors further showcase the benefit of a task-specific approach to automatic fact-checking compared to a suboptimal generalized approach.
+Finally, to understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and identified a heavy skew toward a few highly frequent frames.
+Notably, the frames used in our case studies are among the most frequently evoked frames in PolitiFact fact-checks.
+This alignment highlights the practical relevance of our approach, as it directly addresses the semantic structures most commonly encountered in professional fact-checking.
+color=yellow]This paragraph is almost like after-thought.
+need to present it more naturally and perhaps in better places.
+— doesn’t really feel like there is a better location.
 To summarize, our contributions are as follows:
-  • We proposed a novel paradigm for automatic fact-checking using frame-semantics.
-  • We developed a pilot dataset for automatic fact-checking of real-world claims using large-scale structured data. [^2] https://github.com/anon-naacl25/
-  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, allowing researchers to target high-impact frames for future studies.
-  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^3] https://anon-naacl25.github.io/
-2 Related Works and Background
+  • We proposed a novel paradigm for task-oriented automatic fact-checking using frame-semantics.
+  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data. [^1] https://github.com/anon-sub-repo/fact-checker
+  • We conducted a novel survey of frames evoked in PolitiFact fact-checks (Table 2), enabling targeting high-impact frames for future studies.
+  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo and source code. [^2] https://anon-sub-repo.github.io/
+2 Related Work
+Large-scale datasets.
+Previous studies Aly et al.
+(2021); Gupta et al.
+(2020); Chen et al.
+(2020) often use the term “large-scale datasets” to refer to collections with a large quantity of claims and tables.
+In these cases, the emphasis is on the breadth of the dataset, with a significant number of claims and corresponding tables, typically featuring a modest number of rows per table.
+For instance, FEVEROUS Aly et al.
+(2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tabular structures.
+In contrast, our work addresses high-volume data, where individual tables can contain hundreds of thousands or even millions of rows, as seen in the OECD database used in our case studies, which averages 596,000 rows per table.
+This shift in scale introduces unique challenges for fact-checking systems, particularly in evidence retrieval.
+Automatic fact-checking.
 Recent advances in automatic fact-checking have been largely driven by the integration of large language models (LLMs) and retrieval-augmented generation (RAG) pipelines.
@@
 (2024) enhances fact-checking by using multimodal inputs and iterative reasoning.
-Similarly, FactLLaMA Cheung and Lam (2023a) combines pre-trained LLaMA models with external evidence retrieval to verify claims.
+FactLLaMA Cheung and Lam (2023a) combines a pre-trained LLaMA model with external evidence retrieval to fact-check claims.
 LLM-Augmenter Peng et al.
-(2023) integrates external knowledge sources and providing feedback to improve model accuracy.
-Singhal et al.
-(2024b) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
-For structured data, many approaches have been studied using fine-tuning methods Zhao et al.
+(2023) integrates external knowledge sources and provides feedback to improve model accuracy.
+Singal et al.
+(2024) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
+For structured data, past work has studied using fine-tuning methods Zhao et al.
 (2022); Gu et al.
@@
 Dater Ye et al.
-(2023) is a detailed fact verification system which decomposes claims into sub-questions using LLMs to simplify the fact-checking process and is the best-performing model on the TabFact Chen et al.
-(2020) dataset.
-In this work, we do not focus on fact verification; however, we choose to utilize LLM-based methods for our fact verification step as they do not require retraining the model when new data is used.
-color=green]I felt this statement undermines the significance of our work.
-It focuses on the usage of LLM.
-But the more important essence of our work is the usage of frames.
-2.1 End-to-end Fact-Checking Systems
-ClaimBuster Hassan et al.
-(2017) describes the process of automatic fact-checking as having several core components including a claim matching component, where claims are matched with previous fact checks if they exist, and a knowledge base lookup, which can directly answer questions using internal knowledge, e.g., Wolfram|Alpha.
-Automatic fact-checking systems, like this work, aim to fulfill the role of the knowledge base.
-color=cyan]Is that internal or external?
-More recent studies Guo et al.
-(2022) have decomposed the process into claim detection, evidence retrieval, and verdict prediction and justification.
-color=green]Looks strange to have a subsection for a short paragraph.
-Its connection to the whole section and to the paper’s topic is not clearly explained.
-— paragraph header looks a bit weirder in my opinion.
-should we just remove this subsection?
+(2023) is a fact-verification system which simplifies the fact-checking process by decomposing claims into sub-questions using LLMs.
+It is the best-performing model on the TabFact dataset Chen et al.
+(2020).
+Our work does not focus on fact verification; however, we choose to utilize LLM-based methods for our fact-verification step as they do not require retraining the model when new data is used.
+Applications of frame semantics.
+Madabushi (2024) compares search methods for finding textual entailment facts in RAG-based systems, including a frame-based method.
+The authors found that their frame-based approach significantly improved recall@K compared to other search-based methods.
+Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other domains Neves Ribeiro et al.
+(2023); Gupta et al.
+(2018); Chen et al.
+(2014).
 Figure 1: An example of our proposed paradigm.
@@
 We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment.
-Figure 1 provides an example of how the system processes the statement “Japan’s life expectancy is the 2nd best in Asia” to illustrate this process.
-color=green]Need some better re-organization.
-the first half of this section has several segments, each with a bold-face title.
-Then we suddenly start to have subsections for the 2nd half, where each subsection is just a paragraph.
+Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.” color=yellow]Make it clear in the figure that “Occupy Rank” is the frame.
+Also mention it in the figure caption.
+color=yellow]Why is “life-expectancy” schema listed twice, with slight difference (w/ and w/o measure)?
+color=yellow]In figure 1, spell out FSP as “Frame Semantic Parsing (FSP)” color=yellow]In Figure 1, the yellow boxes are procedures.
+Other colors are data/concept.
+But they are all represented as boxes.
+In addition to color, can we further differentiate them using something else?
+color=yellow]Figure 1 and this section may appear inconsistent.
+“claim understanding”, “evidence retrieval”, “claim-evidence aligenment” is no where to be found in the figure.
+On the other hand, isn’t not clear table mapping, table querying, FSP, fact verfication belong to which of the 3 steps.
+It is like we have presented two alternative naming systems or workflows.
 3.1 Claim Understanding
 In the first step, the system focuses on understanding the specifics of the claim.
-Frame-semantic parsing enables the extraction of structured, predefined representations from claims.
-By identifying frames and their associated frame elements (FEs), the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification.
-For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.
-Frame semantics not only facilitate claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources.
-While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables, other frames may require a predicted mapping based on the text.
-Figure 1 illustrates an example of this predicted mapping.
+Frame-semantic parsing Das et al.
+(2014) enables the extraction of frames and their frame elements from claims.color=green]can make the concept of FSP more clear.
+Even the next sentence (“identifying frames and … FEs …”) kind of defines it and can be used here instead.
+color=green]Need references for FSP By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification.
+For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.color=green]Be consistent.
+The frame name in the figure doesn’t use underscore.
+Frame-semantics not only facilitates claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources.
+While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables (i.e., congress members and bills), other frames may require a predicted mapping based on the text.
+Figure 1 illustrates an example of this predicted mapping.color=yellow]we can find ways to articulate direct vs.
+predicted mapping more clearly.
+Probably won’t have time before deadline.
 3.2 Data Collection and Evidence Retrieval
 The availability of trustworthy data is essential for fact-checking.
-Rather than relying on search engines that can return inconsistent results or be vulnerable to fake news injections Huynh and Hardouin (2023); Horne et al.
-(2019), our system stores ground-truth data in an internal database.
-Using data from reliable sources like the OECD and U.S.
-Congress ensures the integrity of the evidence.
+Our system incorporates hand-selected data from existing reliable sources, such as the OECD and U.S.
+Congress, into an internal database.
+This ensures the integrity of our evidence and avoids the vulnerabilities of external sources.
 Evidence retrieval is guided by the frame elements identified in the claim understanding phase.
@@
 By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy.
-Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by adapting methods like retrieval-augmented generation Lewis et al.
-(2020) using text chunking techniques to query large text corpora instead of a database.
+Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by querying large text corpora instead of a database.
+color=yellow]If we don’t have experiments to support this, we don’t want to discuss it here.
+A future work or discussion section at the end is the right place.
 3.3 Claim-Evidence Alignment
@@
 By only using specific spans of the claim, the system can align the claim and evidence more precisely.
-This approach also enhances explainability, as the system can point to the exact text used in the query.
+This approach also enhances explainability, as the system can point to the text used for evidence retrieval.
 Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim.
 For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best.
-While this claim is not exactly true, the spirit of the claim is true, i.e., that the life expectancy is extremely high in Japan, and thus the model would predict that it is mostly true.color=yellow]This is an intriguing example.
+While this claim is not entirely accurate, its underlying premise holds true—Japan has an exceptionally high life expectancy.
+As a result, the model would predict it as mostly true.
+color=yellow]This is an intriguing example.
 But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it?
@@
 In this section we provide two case studies which utilize our proposed paradigm.
+color=yellow]This is really an important ToDo item. Untify the two case studies in the same terminology framework.
+color=yellow]Another important ToDo item to strengthen the paper: incorporate more case studies.
 4.1 Voting Records
 PolitiFact is a leading fact-checking organization that assesses the accuracy of public statements using verifiable evidence.
-Voting records  [^4] https://www.politifact.com/voting-record/ are a substantial portion of PolitiFact’s fact-checks.
+Voting records are a significant portion of PolitiFact’s fact-checks.
 In this case study, we focus on automatically fact-checking voting-related claims using the Vote frame and official U.S.
@@
 [^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls.
-Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.
-In their fact-checking process, PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims.
-In the context of the Vote frame, claims often reference specific congressional bills.
+Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.color=yellow]We are mentioning 93rd, 101st, and 1st.
+If we only have votes since 101st Congress, why did we include 93rd Congress or even the 1st Congress?
+At least mentioning statistics for them becomes less relevant.
+— this is the data we have, it wouldn’t really make sense to say we only have data from the 101st since we technically have members from before
+PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims.
+444 https://www.politifact.com/article/2022/mar/31/politifacts-checklist-thorough-fact-checking/ In the context of the Vote frame, claims often reference specific congressional bills.
 To evaluate our system, we constructed an evaluation dataset by extracting all PolitiFact fact-checks made before April 2022 that involve claims related to the Vote frame and reference at least one congressional bill.
@@
 To verify voting records, mapping the Agent FE to the correct member of Congress is essential.
-We use SQL queries to match congress members whose names are similar to the words in the Agent FE.
-In cases of name ambiguity, e.g., when two members share the same name, we default to the more recent member.
+We use SQL queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4).
+In cases of name ambiguity we default to the more recent member.
 This stage presents several challenges.
@@
 To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members.
-Although not exhaustive, these lists should cover many common cases.
 Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz.
-We handle this by utilizing a list of congress members’ preferred names, supplemented by common alternatives  [^6] https://github.com/carltonnorthern/nicknames when necessary.
+We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives when necessary.
 Bill matching.
@@
 Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval.
-To address this, we employ an asymmetric semantic similarity model, which allows us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
+To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
 Verifying alignment of claim to bill.
 Determining whether a claim is refuted or supported by evidence presents several challenges.
-First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (FE) (for or against).
+First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (for or against).
+[^6] Note that Position FE is extracted from claims by our frame semantic parsing model.
+However, only Agent and Issue FEs are used for evidence retrieval in voting-related claims.
 Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ.
@@
 Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
-Claim Verification.
+Fact verification.
 Finally, our system integrates the alignments between each bill and claim to perform fact verification over all of the retrieved evidence.
-We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix D.4.
-color=green]Not clear to me what “each individual claim-bill alignement” means.
-Figure 2: System design of OECD case study.
-The frame-semantic parser (orange) facilitates the claim understanding process, table matching and querying (blue) facilitate the evidence retrieval process, and the claim verification step (green) facilitates the fact verification.
-4.2 OECD-Related Statistics
-The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics and serves as a strong focal point to explore a wider range of semantic frames.
-As a visual aid, we include a system overview in Figure 2.
+We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix E.4.
+4.2 OECD Statistics
 Dataset.
-For the OECD Case Study, we collect all of the data tables available on the OECD Data Explorer.
-[^7] https://data-explorer.oecd.org/ We construct a SQLite database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totalling 4.1 billion cells.
+The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics which serve as a strong focal point to explore a wider range of semantic frames.
+For the OECD case study, we collected all of the data tables available on the OECD Data Explorer.
+[^7] https://data-explorer.oecd.org/ We constructed an SQL database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totaling over 4.1 billion cells.
 Relevant table identification.
-To identify relevant tables for data extraction, we apply semantic similarity models to select the five most similar tables based on the extracted frame elements (FEs).
-The query consists of one or more FEs derived from frame-semantic parsing, while the target is a specific table.
-We encode each table’s name and description as text for this purpose.
-Given the brevity of our queries and the length of the table descriptions, we utilize asymmetric semantic similarity models to ensure precise matches.
-To minimize the risk of overlooking relevant data, we retrieve the top five most similar tables.color=green]Is “asymmetric semantic similarity” a research field, a technique, or a particular model (maybe in GPT suite or library)?
-Need to be clear on this and provide link/reference.
-— a type of model determined by how it is trained, i.e., short query vs long document
-Querying Relevant Tables.
-Querying the relevant tables presents challenges, as the system must recognize how values in the claim are represented in the database.
-color=cyan]We said “presents challenges” quite many times in the system.
-You might want to consider add some variety to the wording.
-In cases where direct mappings are available (e.g., in the Vote frame case study), we map FEs to specific tables.
-However, in the OECD case study, where no direct mapping exists between FEs and table columns, the system must predict this mapping.
-To handle these cases, we employ LLM-based code generation, allowing for the execution of multiple simple queries using programmed logic, rather than relying on complex text-to-SQL generation.color=cyan]Should we include some text-to-SQL references?
+To identify the relevant table(s) to fact-check a given claim, we use a RoBERTa-based semantic similarity model Liu et al.
+(2019).
+To represent the table, we encode the table’s name and description as text.
+To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element from the claim.
+Querying retrieved tables.
+Querying the retrieved tables requires understanding of the tables, columns, and the values within them.
+To do this, we encode each table’s name, all of its columns, and several representative sample values from each column into text.
+We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table accordingly.
+We choose this approach, as opposed to standard text-to-SQL approaches because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
 We represent the table schema as SQL code, following best practices from previous work Gao et al.
-(2023), which studied different methods of representing SQL tables for text-to-SQL generation.color=green]What is Gao et al.
-2023 about?
-on the same problem or something else.
-YOu might want to make this clear.
+(2024).
 Alongside the schema, we include representative example values from the database for each column.
-color=green]“Columns ….
-include all values” doesn’t seem correct.
-— referring to how they are encoded into text Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values.
-For columns with a moderate number of unique values, we use a semantic similarity model to select the top ten most relevant values based on the claim.
-Given the large number of cells generated by this process, we use an LLM to refine the query and filter out irrelevant columns, focusing on those most critical for fact-checking the claim.
-The prompt for this LLM can be found in Appendix D.3.
-color=green]Is this semantic similarity model the same as the aymmetric one mentioned in the last section?
-If yes, make it clear; if not, need reference/link, based on the same concern mentioned earlier.
-— each model is different because they perform differently on the data, the selection is discussed in the experiments color=green]Are we missing some details here about how we used LLM?
-Are there prompts?
-Fact Verification
+Our implementation is similar to Nahid and Rafiei (2024); however, we address critical limitations to support the high-volume OECD data.
+Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values.
+For other columns we use the same RoBERTa-based semantic similarity model to select the top ten most relevant values based on the claim.
+These queries often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim.
+The prompts for this process can be found in Appendix E.3.
+Fact verification.
 In the final step, we used an LLM-based fact verification model.
-We represent the extracted evidence as a list of tab-separated values based on the data retrieved in the previous step.
-The model is instructed using the prompt in Appendix D.4 and outputs one of five verdicts: false, mostly false, half-true, mostly true, or true.color=cyan]If time allows, explain the connection to PolitiFact’s Truth-o-meter.
-5 Experiments
+We represent the extracted evidence as a list of tab-separated values using the data retrieved in the previous step.
+The model is instructed using the prompt in Appendix E.4 and outputs one of five verdicts: “false,” “mostly false,” “half-true,” “mostly true,” or “true,” based on PolitiFact’s Truth-o-Meter.
+[^8] https://www.politifact.com/article/2018/feb/12/principles-truth-o-meter-politifacts-methodology-i/
+5 Experiments and Results
 5.1 Datasets
@@
 (1998).
-Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in this dataset along with the newly-defined frames.
-These frames can be found in Appendix 6.
-PolitiFact Fact-checks
-For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles collected as of April 2022.
-Each article contains a claim, a detailed fact-check, a verdict, and a list of sources.
-To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we identified 79 fact-checks that cite congress.gov and evoke the Vote frame.
+Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in their dataset along with the newly-defined frames.
+These frames can be found in Table 6 in the Appendix.
+PolitiFact fact-checks.
+For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles (i.e., fact-checks) collected as of April 2022.
+Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources.
+To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we manually identified 79 fact-checks that cite congress.gov in the sources and evoke the Vote frame.
 For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset.
-color=cyan]PolitiFact fact-checks and PolitiFact fact-check articles.
-Mention very close to each other.
-color=green]the claim cites congress.gov?
-color=green]“bill cited was referenced by the claim”.
-cited where?
-also the claim would reference a bill?
-Additionally, we collected 68 articles that cite oecd.org for our analysis of OECD-related claims.
-Each OECD claim was manually verified and mapped to the statistics which can be used to fact-check it.
-These two sets of claim-evidence pairs—the voting-related claims and the OECD-related claims—serve as evaluation sets for our case studies.
-color=green]these articles cite the website?
-color=green]“To collect … we collect” color=green]Which two sets?
+Similarly, we collected 68 fact-checks that cite oecd.org for our analysis of OECD-related claims.
+Each OECD claim was manually verified and mapped to the database table which can be used to fact-check it.
+Additional details are provided in Appendix A.
 Model Frames Frame Acc FE Acc
@@
 5.2 Frame-Semantic Parsing
-For frame-semantic parsing, we combine the frame identification model developed by Devasier et al.
+For frame-semantic parsing, we combined the frame identification model developed by Devasier et al.
 (2024) with a frame element identification model based on AGED Zheng et al.
 (2023).
-By combining these state-of-the-art approaches, we unify frame and frame element identification into a single model.
-This model is then trained only on the fact checking frames introduced in Arslan et al.
+By combining these state-of-the-art approaches, we unified frame and frame element identification into a single RoBERTa-based model (additional details in Appendix C.4).
+We fine-tuned two of these frame-semantic parsing models, one for each case study, using their respective frames from the training data in Section 5.1.
+The model used by our system is then selected based on the candidate targets in the sentence (extracted using the approach in Devasier et al., 2024).
+To evaluate the performance of our model for the case studies, we compared it with a baseline GPT-4o-mini model using OpenAI’s structured generation.
+The prompt used for this model can be found in Appendix E.1.
+We evaluated each model using exact match accuracy on frame identification and argument identification.
+These results are presented in Table 1.
+Our findings are consistent with a previous study indicating that generative LLMs tend to struggle with frame-semantic parsing Su et al.
 (2024).
-To evaluate the performance of our model for the case studies, we compare it with a GPT-4o-mini model using OpenAI’s structured generation [^8] At the time of submission, GPT-4o-mini is the only model which supports structured generation..
-We evaluate each model using accuracy on frame identification and exact match accuracy on FE identification in Table 1.
-We define a separate prompt for each case study in Appendix D.1.
-5.3 Claim Coverage
-One concern which may arise is regarding the applicability of our system for assisting in the fact checking process.
-To understand this, we aim to identify the coverage of a frame-based approach towards automatically fact-checking claims.
-Due to the lack of annotations for all of the fact-checking frames in Arslan et al.
-(2020), we utilize a zero-shot GPT-4o-mini model to identify the frames evoked in the claim of each PolitiFact fact-check (Table 2).
-Claims which do not evoke a frame studied in this work are categorize as none.
-The prompt used for this model can be found in Appendix D.1.
+Our fine-tuned frame-semantic parsing model performed much better on frame and argument identification for both case studies.
+A key observation was GPT-4o-mini’s tendency to over-predict frame occurrence in sentences.
+Performance improved significantly when we constrained the analysis to only the first predicted frame (GPT-4o-mini*).
+Similar patterns emerged in frame element predictions, where GPT-4o-mini frequently predicted frame elements that do not exist in the frame definition or are not used in the input sentence.
+Examples of these are provided in Appendix D.2.
+5.3 PolitiFact Survey
+To assess the broader applicability of our frame-semantics-based fact-checking system, we conducted a comprehensive coverage analysis.
+Because the annotations available for the fact-checking frames are limited, we employed a zero-shot GPT-4o-mini model (detailed in Appendix E.1) to identify frames evoked in the PolitiFact fact-check corpus.
+Claims that did not trigger any of our studied frames were classified as “None.” The distribution of frames is presented in Table 2.
+Our findings show that GPT-4o-mini may predict more frames than there actually are.
+This is because the annotated dataset evokes an average of 1.1 frames per claim while GPT-4o-mini predicted an average of 2.1 frames per claim.
+We believe this is because GPT-4o-mini predicts frames for lexical units–words which evoke frames–that are not defined by Arslan et al.
+(2020).color=yellow]This last sentence appears to suggest GPT can find frames there were missed in Fatma’s dataset.
+But why did you say it “predict more than there actually are”?
+— This isn’t really known 100% the cause could be that the model over-predicts in a bad way or it could be because the model recognizes frames for more LUs than are defined.
+It seems to be the latter based on what ive seen.
+One indication of this can be seen with the Vote frame.
+While it explicitly defines only the lexical unit vote.v, the frame could be evoked through various other expressions.
+For instance, the statement “I passed a bill” implicitly suggests a vote of affirmation.
+Our analysis found that only 1,198 (5.7%) of PolitiFact fact-checks explicitly mention “vote,” whereas the GPT-4o-mini model predicted Vote frames at nearly three times that amount.
+Although we observed numerous claims that evoked the Vote frame without explicitly using the word “vote,” additional research is needed to determine the true distribution of frames in factual claims.
 Model Data R@K
+text-embedding-3-large Vote 0.032
+BM25 Vote 0.048
+roberta-base-v2 Vote 0.131
+stella_en_1.5B_v5 Vote 0.144
+distilbert-tas-b Vote 0.165
 BM25 OECD 0.323
@@
 roberta-base-v2 OECD 0.726
-text-embedding-3-large Vote 0.032
-stella_en_1.5B_v5 Vote 0.144
-distilbert-tas-b Vote 0.165
 Table 3: Performance comparison of different semantic similarity models on extracted frame elements. K=10 for Vote and K=5 for OECD.
+5.4 Evidence Retrieval
+Similarity model selection.
+To retrieve relevant information for fact-checking, we first evaluated different semantic similarity models.
+Table 3 presents the embedding models we compared, including Stella Zhang et al.
+(2025), RoBERTa Liu et al.
+(2019), DistilBERT-TAS-B Hofstätter et al.
+(2021), OpenAI’s text-embedding-3, 999 https://platform.openai.com/docs/models and BM25.
+101010 https://github.com/dorianbrown/rank_bm25 color=yellow]In Table 4, font size of different models are different; some R@K is italic, others not.
+Table 1 and Table 3/4 use different font styles.
+Our analysis revealed that different models excel in specific verification contexts.
+RoBERTa demonstrated superior performance in identifying relevant OECD tables, while DistilBERT-TAS-B achieved the best results in matching claims to bills.
+These findings suggest that optimal fact-checking systems should employ task-specific embedding models rather than a one-size-fits-all solution.
 Model Query Data R@K
+distilbert-tas-b Full claim Vote 0.143
+distilbert-tas-b Issue FE Vote 0.165
+Max Possible - Vote 0.568
 RoBERTa (v2) Full claim OECD 0.653
@@
 Max Possible - OECD 0.910
-distilbert-tas-b Full claim Vote 0.143
-distilbert-tas-b Issue FE Vote 0.165
-Max Possible - Vote 0.568
-Table 4: Performance of different methods of representing the query in the OECD case study’s table matching. OECD and Vote claims use Recall@5/10, respectively.
-5.4 Evidence Retrieval
-We study the capability of our system to retrieve the relevant information which can be used to fact-check the claim.
-To do this, we found the percentage of bills and statistics available in our database for voting and OECD-related claims, respectively.
-This serves as a performance ceiling on our collected claims as our system would be unable to fact-check claims which do not have available data.
-These values are found in Table 4 (Max Possible).
-For semantic similarity, we compare commonly used embedding models (Table 3), namely RoBERTa Liu et al.
-(2019) and DistilBERT-TAS-B Hofstätter et al.
-(2021).
-Additionally, we compare with OpenAI’s text-embedding-3-large model and provide a baseline BM25 approach.
-We also experiment with different approaches of representing the query used to find relevant data.
-In Table 4, we compare the performance of using the entire claim with only the relevant FE.
-To evaluate the performance of these queries, we calculate the recall@K, where K=5 for OECD claims and K=10 for Vote claims.
-The relevant FEs used for each frame can be found in Appendix A.
+Table 4: Performance of different query representations in each case study. Evaluations are performed on table matching (OECD) and bill matching (Vote), and use Recall@5/10, respectively.
+Query representation.
+We explored two different methods for constructing retrieval queries: using the complete claim text and using only the frame elements (FEs) extracted based on the evoked frame (detailed in Appendix B).
+We evaluated these query representations using the best-performing similarity models identified in Table 3.
+Table 4 presents the recall@K metrics for each representation approach, with K=5 for OECD claims and K=10 for Vote claims.
+We also established baseline metrics for evidence availability (Max Possible) in our database.
+For voting-related claims, we calculated the percentage of referenced bills present in our database, while for OECD claims, we assessed the availability of relevant datasets.
+Our results demonstrate that queries constructed from frame elements consistently outperformed full-claim queries across both domains.
+This finding aligns with research by Madabushi (2024) showing that frame-semantics-based search strategies enhance retrieval effectiveness.
+The performance varied significantly between domains.
+For voting-related claims, even our best-performing model (DistilBERT-TAS-B) achieved a modest recall@10 of 0.165, reflecting the complexity of bill matching.
+In contrast, OECD claim verification was easier for the model, with RoBERTa achieving a recall@5 of 0.726, likely due to the more constrained search space.
+Congress member identification performed very well with an exact match accuracy of 97.5%.
+Query generation.
+The effectiveness of generated queries varied significantly between the two domains: 62% of OECD claim queries successfully retrieved relevant data, compared to only 36% for voting claims.
+These retrieval rates effectively establish upper bounds for our fact verification capabilities, as meaningful fact-checking requires successful evidence retrieval.
 Model Dataset Accuracy
 GPT-4o Naive Vote 0.044
-Our w/ Irrelevant Vote 0.076
-Our w/o Irrelevant Vote 0.207
-Our w/ Irrelevant ± Vote 0.253
-GPT-4o Naive ± Vote 0.324
-Our w/o Irrelevant ± Vote 0.690
+Ours w/ Irrelevant Vote 0.076
+Ours w/o Irrelevant Vote 0.207
 GPT-4o Naive OECD 0.073
-Our w/ Irrelevant OECD 0.214
-Our w/o Irrelevant OECD 0.429
-GPT-4o Naive ± OECD 0.537
-Our w/ Irrelevant ± OECD 0.607
-Our w/o Irrelevant ± OECD 0.821
-Table 5: Fact verification performance on different case studies.
-± indicates the off-by-one performance.
-Bolded values indicate the best-performing settings overall and underlined values are only considering exact matches.
+Ours w/ Irrelevant OECD 0.214
+Ours w/o Irrelevant OECD 0.429
+Table 5: End-to-end fact verification performance on the two case studies. The best-performing models are bolded.
 5.5 Fact Verification
-Finally, we study the ability of our fact verification system to fact-check claims using structured data by comparing the verdicts presented in each PolitiFact fact-check to the model’s predicted verdict.
-We have replaced PolitiFact’s “pants on fire” verdicts with a false verdict.
-To determine whether the LLM is relying on internal knowledge or if it is using the provided data, we evaluate the performance of GPT-4o without any data.
-We measure accuracy as the primary evaluation metric and also calculate off-by-one accuracy, where, for example, a prediction of “mostly false” is considered correct if the label is “false.” The results of these experiments are summarized in Table 5.
-6 Results
-6.1 Frame-Semantic Parsing
-We evaluated the frame-semantic parsing model on fact-checking frames introduced by Arslan et al.
-(2020), comparing it against GPT-4o-mini due to its structured output capability.
-Table 1 shows the performance on both the Vote and OECD frames.
-We found that the GPT-4o-mini model tends to over-predict the number of frames in a sentence.
-When we only use the first predicted frame, the performance is increased significantly.
-We saw similar cases with frame element predictions where GPT-4o-mini tends to predict more frame elements than actually exist in the input.
-We include examples of each of these cases in Appendix C.2.
-6.2 PolitiFact Survey
-Similar to Section 6.1, we found that GPT-4o-mini tends to predict more frames than there may actually be.
-We believe this is likely due to the annotations from Arslan et al.
-(2020) being sampled from PolitiFact and likely having a similar structure.
-We found the annotated samples evoked 1.1 frames on average while GPT-4o-mini predicts an average of 2.1 frames.
-Another possible explanation is that the model is able to identify frames for lexical units– words which evoke frames–which are not defined.
-One indication of this is on the Vote frame, which only has a lexical unit defined for vote.v; however, it is possible to evoke the Vote frame using other words.
-For example, the statement “I passed a bill” implies that a vote of affirmation was cast on that bill.
-We found only 1,198 (5.7%) PolitiFact fact-checks which mention “vote” while the model predicts that almost 3 times as many evoke the Vote frame.
-While we found many claims which evoke a Vote frame without an explicit mention of the word “vote”, further analysis is needed to get a more accurate understanding of the true distribution of frames evoked by factual claims.
-6.3 Evidence Retrieval
-We found that embedding models’ performance is task-specific and there is no single-best embedding model for every task.
-For example, RoBERTa performed the best for finding the right OECD table while DistilBERT-TAS-B was better for matching claims to their corresponding bills.
-We also found that, in general, using the claim-specific frame element performs better than using the entire claim as a query, as shown in Table 4.
-For voting-related claims, we found that identifying the bill referenced in a particular claim is very difficult.
-Our best performing model, DistilBERT-TAS-B, was only able to obtain a Recall@10 of 0.165.
-For OECD-related claims, identifying the table to use to fact-check the given claim was much easier, with the best model, RoBERTa, achieving a recall@5 of 0.726.
-This is likely because the similarity between the relevant FE and the table names were better-represented by semantic similarity models.
-This indicates a potential area of significant improvement in the methodology and representation of bill texts according to their topics and implications, in addition to the text itself.
-6.4 Fact Verification
-On OECD claims, 62% of the generated queries were able to retrieve relevant data while only 36% of the voting claims were able to retrieve relevant bills.
-These values put an upper bound on the capabilities of our fact verification model as without evidence, the model will not be able to provide useful fact-checks.
-In general, we found that language models tend not to predict a claim to be true based on the given data, yet they have no issue with predicting a claim to be false, and often do so.
-We believe this is analogous to proofs, it is quite easy to prove by contradiction, if you find a single vote or country statistic which contradicts the claim, while it is much harder to assert that a claim is true without looking at all of the data.
-We found that on Vote claims, a naive GPT-4o correctly predicted the truthfulness of the factual claim with an accuracy of 0.286; however, when voting data is added, the performance drops to 0.076.
-This suggests that, while there may be some internal knowledge on certain people and bills, it is unlikely that the model is relying on it when given specific voting records.
-This also indicates that the prompt given to the model is sufficient in preventing the LLM from relying on internal knowledge.
-When excluding claims for which data was not found, we observed absolute performance increases of +13.1% and +21.5% for Vote and OECD claims, respectively.
-We also found that, while the exact-match accuracy of the system is low, there is a strong correlation between predictions and labels, as indicated by the off-by-one accuracy in Table 5.
-We include example outputs for the explanations of these models in Appendix C.1.
+We evaluated our fact verification system by comparing its predicted verdicts against the ground truth verdicts from PolitiFact fact-checks (Table 5).
+For consistency in our analysis, we consolidated PolitiFact’s "pants on fire" rating with "false".
+To assess whether our system genuinely relies on retrieved evidence rather than pre-existing knowledge, we established a baseline using GPT-4o without access to any external data (GPT-4o Naive).
+Our analysis encompassed claims across both voting records and OECD datasets, with performance measured against human-annotated verdicts.
+The baseline GPT-4o Naive model performed poorly on both case studies, suggesting that the model’s internal knowledge alone is insufficient for accurate fact verification.
+When we excluded cases where our system failed to retrieve relevant evidence (w/o Irrelevant), we observed substantial improvements in accuracy: an increase of 13.1 percentage points for voting-related claims and 21.5 percentage points for OECD-related claims.
+This improvement demonstrates the critical role of successful evidence retrieval in fact verification accuracy.
+Detailed examples of the system’s reasoning and explanations can be found in Appendix D.1.
+Our analysis revealed a bias in the model’s predictions toward classifying claims as false.
+This bias likely stems from the nature of contradiction-based verification where disproving a claim requires finding just one contradictory piece of evidence, while proving a claim often necessitates comprehensive evidence supporting all aspects of the assertion.
+This suggests potential areas for future improvement in balancing verification criteria.
 6 Conclusion
-This research contributes to the evolving field of automatic fact-checking by proposing a novel paradigm that incorporates frame semantics to enhance the evidence retrieval process and enable fine-grained control of downstream tasks.
-By developing and presenting a pilot dataset specifically annotated for real-world claims, this work helps to address the underexplored challenge of fact-checking using massive, structured data sources.
-Our findings demonstrate the promise of frame-semantic parsing in improving evidence retrieval for automatic fact-checking.
-In both case studies, we observed that using frame elements extracted from claims instead of entire claims led to substantial improvements in recall metrics, highlighting the utility of frame-based understanding in identifying relevant evidence.
-These results suggest that semantic frames not only offer a structured and explainable approach to parsing claims but also guide retrieval models more effectively toward the correct information within large datasets.
-Looking ahead, the integration of frame semantics into fact-checking systems holds promise for further advancements in both accuracy and interpretability.
-Future work could explore broader applications of this approach, including expanding the dataset, refining the models, and incorporating additional semantic frames to cover more claims.
+In this work, we introduced a novel paradigm for task-oriented automatic fact-checking by leveraging frame semantics to enhance structured understanding and evidence retrieval.
+Through the development of a pilot dataset composed of real-world claims and two focused case studies, we demonstrated how frame-semantic parsing can guide the fact-checking process with structured data, improving retrieval performance and explainability.
+Our findings suggest that semantic frames offer a meaningful way to bridge the gap between complex claims and high-volume structured datasets.
+The empirical results highlight the advantages of using frame elements for evidence retrieval, with improvements in recall across both voting-related and OECD-based fact-checking tasks.
+By aligning claims with relevant structured data, our approach not only enhances accuracy but also provides a more interpretable verification process.
+Furthermore, our survey of fact-checking frames reveals promising avenues for future research, particularly in expanding the coverage of high-impact frames and refining claim-evidence alignment techniques.
+While our system is limited by the availability of trustworthy data, our findings highlight a critical research gap left by approaches that focus on smaller tabular data.
+This work showcases the use of high-volume structured databases for automatically verifying novel claims.
+Expanding beyond our pilot datasets to large-scale collections of claims and corresponding evidence from structured sources remains an important next step.
 Limitations
 Despite the advancements made in this study, our system’s limitations should be acknowledged.
-While this study focuses primarily on improving evidence retrieval and frame-semantic parsing, it does not comprehensively evaluate different fact verification approaches.
+While this study focuses primarily on improving evidence retrieval, it does not comprehensively evaluate different fact verification approaches.
 Fact verification is critical for automatic fact-checking, and a deeper investigation into various verification strategies, including fine-tuned models and LLM-based approaches, could further enhance system performance.
@@
 Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
-Appendix A Studied Frames
-Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings.
+Appendix A Data Quality
 Frame Frame Element Frame Definition
@@
 Capability Event An Entity meets the pre-conditions for participating in an Event.
+Vote Agent, Issue An Agent makes a voting decision on an Issue.
 Table 6: Frames used for OECD case study, along with the frame elements extracted from corresponding frames.
-Appendix B Reproducibility
-We trained our frame-semantic parser using a NVIDIA GTX 1080 Ti in about 6 hours on a single run without any hyperparameter tuning. Github Copilot was used in the development of some of our code.
-B.1 Software License
-Our source code and data are released with the MIT License.
+Annotators.
+The OECD and voting-related claims were annotated by three students, including one senior PhD student with expertise in frame-semantic parsing.
+The PhD student provided the final ground truth annotations.
+The two other students independently provided their annotations, and if they matched the PhD student’s annotations, they were kept.
+No disagreements occurred during annotation.
 Model Parameters URL
@@
 text-embedding-3-large unknown text-embedding-3-large
-Table 7: Models and Their Corresponding Hugging Face URLs
-B.2 Models
-Unless stated otherwise, all models used are the “base” variant.
-All semantic similarity models are available via the sentence-transformers library.
-[^9] https://huggingface.co/sentence-transformers Table 7 contains all of the models used and their details.
-Claim Predicted Verdict Explanation Fact Check
+Table 7: Models and their corresponding Hugging Face URLs.
+Voting claims.
+Our dataset of voting-related claims consists of original claims collected from PolitiFact, corresponding URLs, manually cleaned claim text, and references to legislative bills cited in the sources.
+The dataset contains 79 claims referencing 190 bills, averaging 2.4 bills per claim.
+However, this distribution is skewed, with a median of only 1 bill per claim due to a few claims referencing a large number of bills.
+Fact-checking activity was concentrated in election years, with 86% of the 79 claims fact-checked in 2016, 2018, or 2020.
+Health insurance was the most frequently mentioned topic, appearing in 11 claims.
+The terms “for” and “against” were used in 28 and 27 claims, respectively, suggesting a frequent focus on arguments in favor of or opposition to specific policies.
+No single individual was disproportionately fact-checked among the 79 claims; Donald Trump had the most fact-checked claims (3), followed by Hillary Clinton (2).
+Claim lengths ranged from 7 to 62 words, with an average of 18.4 and a median of 17 words.
+OECD claims.
+The OECD-related claims were more evenly distributed across years, with the top three years (2016, 2014, 2011) accounting for 46% of the 68 claims.
+The United States was the most frequently mentioned country, appearing explicitly in 25% of the claims.
+Given PolitiFact’s U.S.
+focus, implicit references to the U.S.
+are likely higher.
+Taxes were the most common topic, with approximately 20% of the claims mentioning tax-related topics. Claim lengths ranged from 9 to 41 words, with an average of 17.7 and a median of 15 words.
+Almost all OECD claims contained a single frame, with only one instance containing multiple frames (Occupy_rank and Comparing_two_entities).
+The most frequent frames were Comparing_two_entities (25 claims), Occupy_rank_via_superlatives (12 claims), and Occupy_rank (8 claims), indicating a strong emphasis on ranking and comparison within these claims.
+Claim Bill Title Vote Alignment Prediction
+Marco Rubio voted against the bipartisan Violence Against Women Act 112 S 1925 Violence Against Women Reauthorization Act of 2012 Nay Supports True
+17 HR 3233 National Commission to Investigate the January 6 Attack on the United States Capitol Complex Act Nay Irrelevant
+117 HR 350 Domestic Terrorism Prevention Act of 2022 Nay Irrelevant
+Chuck Grassley was voting to slash Medicare when voting against the debt ceiling bill 117 S 610 Protecting Medicare and American Farmers from Sequester Cuts Act Nay Supports Mostly False
+117 S 1301 Promoting Physical Activity for Americans Act Nay Irrelevant
+17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
+Table 8: Examples of voting-related claims with the corresponding retrieved bills, votes on the bills, vote-claim alignment, and fact verification prediction.
+Appendix B Studied Frames
+Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings.
+The selected frames were chosen based on their relevance to factual claims in the domains of voting and country statistics.
+Our analysis primarily focused on frames that involve ranking, comparisons, and legislative actions, as these are commonly invoked in fact-checking scenarios.
+Appendix C Reproducibility
+Ensuring reproducibility is a core principle of this work. We provide details about our experimental setup, computational resources, software dependencies, and models used.
+C.1 Computational Resources
+We trained our frame-semantic parser on an NVIDIA GTX 1080 Ti GPU.
+A single training run took approximately 6 hours.
+No hyperparameter tuning was performed, and the model was trained using default settings unless explicitly stated.
+The OECD database requires roughly 85GB of storage and the voting records requires 1GB of storage.
+color=cyan]we don’t train any other models.
+C.2 Software and Dependencies
+Our implementation is based on PyTorch and Hugging Face’s “transformers” and “sentence-transformers” libraries.
+A full list of dependencies and their versions is provided in our public repository.
+GitHub Copilot was used in the development of some of our code to assist with boilerplate functions and standard NLP preprocessing tasks.
+C.3 Software License
+All source code and dataset annotations are released under the MIT License.
+C.4 Models
+Unless stated otherwise, all models used in our experiments are the “base” variant.
+All semantic similarity models are sourced from the “sentence-transformers” library.
+[^11] https://huggingface.co/sentence-transformers Table 7 provides an overview of the models used and their parameter count.
+Frame-semantic parsing.
+Our frame-semantic parsing model is an adaptation of AGED, augmented with the frame classification layer and strategy from Devasier et al.
+(2024).
+We fine-tuned this model using a joint loss function to improve frame prediction accuracy.
+Instead of BERT, as originally used in AGED, we employ RoBERTa for its improved contextual embeddings, which have been shown to enhance performance on frame-based semantic parsing tasks.
+Voting agent matching.
+Our agent matching function retrieves the BioGuide ID of a political figure from a database.
+It extracts the agent’s name using frame element spans (FEs) and checks if it is a known presidential nickname, replacing it with the full name if applicable.
+If not, it uses a SpaCy NER model to extract PERSON named entities.
+The function first attempts an exact match in the members table, then a fuzzy match using SQL’s LIKE operator.
+If neither succeeds, it constructs a multi-condition query, incorporating nickname expansions where applicable.
+Appendix D Examples
+To provide a clearer understanding of our system’s outputs, we include example cases from both the voting and OECD datasets.
+These examples illustrate how our models process claims, align them with retrieved documents, and generate predictions.
+Claim Predicted Verdict Explanation Fact-Check
 Americans work way more than an average of industrialized countries around the world.
@@
 PolitiFact
-Table 8: OECD fact check example predictions and explanations
-Appendix C Examples
-In this section, we provide a few example outputs of our system to give a better understanding of the outputs of our models.
-Claim Bill Title Vote Alignment Prediction
-Marco Rubio voted against the bipartisan Violence Against Women Act 112 S 1925 Violence Against Women Reauthorization Act of 2012 Nay Supports True
-17 HR 3233 National Commission to Investigate the January 6 Attack on the United States Capitol Complex Act Nay Irrelevant
-117 HR 350 Domestic Terrorism Prevention Act of 2022 Nay Irrelevant
-Chuck Grassley was voting to slash Medicare when voting against the debt ceiling bill 117 S 610 Protecting Medicare and American Farmers from Sequester Cuts Act Nay Supports Mostly False
-117 S 1301 Promoting Physical Activity for Americans Act Nay Irrelevant
-17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
-Table 9: Examples of voting-related claims with the corresponding retrieved bills, votes on the bill, the vote-claim alignment, and fact verification prediction.
-C.1 Vote-Claim Alignment
-We include two sample claims and their corresponding top three retrieved bills along with the vote-claim alignment for each bill in Table 9.
-C.2 Fact Verification Predictions
-Vote
-We include two claims and their corresponding predictions based on the retrieved bills in Table 9.
-OECD
-We include two predictions and their explanations on OECD claims in Table 8.
-Appendix D LLM Prompts
-D.1 Frame-Semantic Parsing
+Table 9: OECD fact-check example predictions and explanations.
+D.1 Vote-Claim Alignment
+Table 8 presents two sample claims from the voting dataset along with their top three retrieved bills.
+For each retrieved bill, we include the computed vote-claim alignment score, which indicates the relevance of the bill to the claim.
+This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.color=red]I don’t any score in the table.
+D.2 Fact Verification Predictions
+Voting claims.
+For voting-related claims, we provide two examples where our model predicts the veracity of claims based on retrieved bills.
+Table 8 includes the claims, retrieved evidence, and the final prediction score.color=red]no score
+OECD claims.
+For OECD-related claims, we provide two sample predictions and their corresponding explanations in Table 9.
+These examples demonstrate how our model interprets numerical comparisons, tax-related claims, and ranking statements to verify factual assertions.
+Appendix E LLM Prompts
+E.1 Frame-Semantic Parsing
 Listing 1: Frame-semantic parser for the Vote frame.
@@
 - Issue: The subject or matter that the Agent is voting on with a particular position.
-- Side: Additional person(s) involved in the voting on the same Issue alongside the Agent.
-- Position: The stance the Agent takes, expressing whether their vote is in favor or against the Issue.
-- Frequency: How often the Agent has made this voting decision regarding the Issue.
-- Time: When the Agent performed the voting decision.
-- Place: Location where the vote took place.
-- Support_rate: The percentage of the Agents total votes aligning with the Side over a set period.
+... [omitted for brevity]
 # Notes
@@
 - Consider synonyms or variations of terms related to voting when analyzing the sentence.
-- Frame elements should quote exactly from the input
+- Frame elements should quote exactly from the input.
 Listing 2: Frame-semantic parser for OECD frames.
@@
 - Rank: Rank held, often defined by a superlative.
-- Dimension: Aspect along which the ranking occurs (e.g., speed, age).
-- Comparison_set: Group of entities being compared.
-- Time: Time period when the item occupies the rank.
+... [omitted for brevity]
 - Occupy_rank:
@@
 - Rank: Rank held.
-- Dimension: Criterion used for ranking.
-- Comparison_set: Set of entities being ranked.
-- Time: Time period when the item holds the rank.
+... [omitted for brevity]
 - Change_position_on_a_scale:
@@
 - Attribute: Property or scale of the change.
-- Difference: Amount of change.
-- Final_value: Final position on the scale.
-- Initial_value: Initial position on the scale.
-- Path: Progression between positions on the scale.
-- Speed: Rate of change.
-- Correlated_variable: Related variable that changes with the Attribute.
-- Manner: Method of performing the action.
-- Degree: Extent to which the change occurs.
-- Circumstances: Context or conditions for the change.
-- Result: Outcome of the change.
-- Group: Collection of Items undergoing change.
-- Time: Time frame during which the change occurs.
-- Duration: Length of time for the change to happen.
-- Value_range: Range of values along the scale.
-- Initial_correlate: State corresponding to the Initial_value.
-- Final_correlate: State corresponding to the Final_value.
-- Place: Location where the Attribute is measured.
-- Initial_state: State of the Item before the change.
-- Final_state: State of the Item after the change.
-- Period_of_iterations: Duration from start to stop of repeated changes.
-- Particular_iteration: Specific instance of a repeated event.
-- Containing_event: Broader event encompassing the change.
-- Explanation: Rationale for the change.
+... [omitted for brevity]
 - Comparing_two_entities:
@@
 - Entity_2: Entity serving as the comparison point.
-- Comparison_criterion: Criterion for comparison.
-- Degree: Extent of the comparison.
-- Time: Period over which the comparison takes place.
+... [omitted for brevity]
 - Comparing_at_two_different_ points_in_time:
@@
 - First_point_in_time: First time period in comparison.
-- Second_point_in_time: Second time period in comparison.
-- Comparison_criterion: Criterion being compared.
-- Degree: Extent of the difference.
-- Difference: Magnitude of change between the two times.
+... [omitted for brevity]
 # Steps
@@
 - If a frame element does not explicitly appear in the claim, leave it blank.
-D.2 Claim-Bill Alignment
+E.2 Claim-Bill Alignment
 Listing 3: Claim-Bill Alignment
@@
 Irrelevant - The vote on this bill is not relevant to the claim at all.
-D.3 OECD Data Query
+E.3 OECD Data Query
 Listing 4: Retrieve Data for Fact-Checking Claim
@@
 - Add more filters to narrow down to claim-relevant metrics; try to have one value per non-numeric column wherever possible.
-- Claim-relevant metrics may appear in columns other than ‘’measure’‘, so reference the claim to identify the appropriate columns to filter by.
+- Claim-relevant metrics may appear in columns other than measure’, so reference the claim to identify the appropriate columns to filter by.
 - Use the retrieval code to identify useful information for your filters, if needed.
-- When filtering by ‘’unit_of_measure’‘, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
+- When filtering by unit_of_measure’, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
 Output:
-- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return ‘’Data is not Available’‘.
-D.4 Fact Verification
+- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return Data is not Available’.
+E.4 Fact Verification
 Listing 6: Vote - Verify Claim and Provide Verdict

### v3

@@
 Abstract
-We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims and guide the process of fact-checking them.
-To support this, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data.
-This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames based on data sources from the Organisation for Economic Co-operation and Development (OECD).
-Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval and explainability for fact-checking.
-Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future work in this direction.
+Automatic fact-checking is a critical tool for addressing the growing challenge of misinformation, particularly when verifying novel claims that lack previously curated evidence.
+Prior work has largely focused on unstructured or highly-curated data sources, limiting scalability and generalization.
+In this work, we introduce a novel paradigm for task-oriented automatic fact-checking using frame semantics to improve interpretability and evidence retrieval from high-volume structured datasets.
+We present a new pilot dataset of real-world factual claims grounded in two large-scale databases: U.S.
+congressional voting records and OECD country statistics.
+Across two case studies using these datasets, we demonstrate how frame elements can guide fine-grained retrieval and automatic fact-checking.
+Our experiments show that frame element-driven evidence retrieval improves recall by 14% and 11% over full-claim baselines in the voting and OECD case studies, respectively.
+We further analyze frame distributions in PolitiFact fact-checked claims and find a strong alignment with the frames targeted in our study.
+Overall, our results highlight frame semantics as a promising foundation for scalable, interpretable, and domain-adaptable automatic fact-checking.
 Task-Oriented Automatic Fact-Checking with Frame-Semantics
@@
 To bridge this gap our work proposes a novel pilot dataset of real-world factual claims related to two high-volume structured databases on U.S.
-congressional voting records and country statistics.color=yellow]Saying we created the dataset to fill the gap gives me the feeling of we invent the nail.
-Instead, consider saying the verification of many claims rely on data that only exist in much larger tables.
+congressional voting records and country statistics.
 These databases significantly differ from past works in both volume and detail of their tables.
@@
 (2020) while our database of country statistics has an average of 596,000 rows across over 400 tables.
-This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed, even for those solutions designed for large tables Nahid and Rafiei (2024).
+This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed due to context length limitations, even for those solutions designed for large tables Nahid and Rafiei (2024).
 Fact-checking novel claims requires fine-grained methodology to understand the claim and to map important parts of the claim to their relevant data.
@@
 Frame semantics is a linguistic framework that explores how meaning is encoded through structured representations called frames.
-These frames capture the important elements and relationships in various situations and enable explainable understanding of claims and extraction of their key parts.
-The existence of specific frames in a claim also enables fine-grained task-specific handling of claims for fact-checking.
+These frames capture the essential elements and relationships of the situations described by claims, supporting explainable understanding and extraction of key components from the claims.
+The existence of specific frames in a claim also enables fine-grained task-specific handling of the claim for fact-checking.
+In our paradigm, we use the frames evoked, and manually-selected frame elements, in a given claim to identify relevant database tables for automatically fact-checking it.
 Our proposed pilot dataset serves as the foundation for two case studies that demonstrate how frame semantics can be used to guide specific fact-checking processes and enhance the explainability of automatic fact-checking.
@@
 These Issue and Agent elements are then used to implement voting-specific fact-checking methods.
-The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
+The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, highly diverse structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
 To evaluate these case studies, we collected 79 and 68 real-world claims for the voting records and OECD statistics, respectively.
@@
 This alignment highlights the practical relevance of our approach, as it directly addresses the semantic structures most commonly encountered in professional fact-checking.
-color=yellow]This paragraph is almost like after-thought.
-need to present it more naturally and perhaps in better places.
-— doesn’t really feel like there is a better location.
 To summarize, our contributions are as follows:
   • We proposed a novel paradigm for task-oriented automatic fact-checking using frame-semantics.
-  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data. [^1] https://github.com/anon-sub-repo/fact-checker
-  • We conducted a novel survey of frames evoked in PolitiFact fact-checks (Table 2), enabling targeting high-impact frames for future studies.
-  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo and source code. [^2] https://anon-sub-repo.github.io/
+  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data and released it along with our source code.
+[^1] https://github.com/idirlab/claimlens-case-studies
+  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, enabling targeting high-impact frames for future studies.
+  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^2] https://idir.uta.edu/claimlens/
 2 Related Work
@@
 For instance, FEVEROUS Aly et al.
-(2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tabular structures.
-In contrast, our work addresses high-volume data, where individual tables can contain hundreds of thousands or even millions of rows, as seen in the OECD database used in our case studies, which averages 596,000 rows per table.
-This shift in scale introduces unique challenges for fact-checking systems, particularly in evidence retrieval.
+(2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tables.
+However, a key limitation of these datasets is their reliance on curated data, largely extracted from Wikipedia.
+Wikipedia tables, while useful for research, often represent a simplified or filtered version of underlying data from primary sources.
+This includes implicit decisions about which metrics to include (e.g., GDP per capita in USD vs.
+PPP), how to aggregate regional data, or how to normalize column formats for readability.
+In contrast, our work avoids these curated intermediaries and instead operates directly on primary source data in its original, often more complex format.
+For example, the OECD tables used in our study contain nearly 600,000 rows per table on average.
+This shift from curated to raw primary data introduces new challenges, particularly in evidence retrieval and reasoning, as systems must handle large-scale, high-dimensional tables without relying on pre-filtered or human-edited summaries.
+It also better reflects real-world fact-checking settings, where claims must be verified against the same raw data sources used by policymakers or researchers.
 Automatic fact-checking.
@@
 (2020).
-Our work does not focus on fact verification; however, we choose to utilize LLM-based methods for our fact-verification step as they do not require retraining the model when new data is used.
+Our work does not primarily focus on fact verification; rather, our main emphasis is on evidence retrieval.
+Nevertheless, we employ LLM-based methods for the fact-verification step, as they offer the practical advantage of not requiring model retraining when new data becomes available.
 Applications of frame semantics.
@@
 The authors found that their frame-based approach significantly improved recall@K compared to other search-based methods.
-Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other domains Neves Ribeiro et al.
+Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other NLP domains Neves Ribeiro et al.
 (2023); Gupta et al.
@@
 Figure 1: An example of our proposed paradigm.
-First, the frame and frame elements (FEs) are extracted using a frame-semantic parser (FSP).
+First, the frame Occupy Rank and frame elements (FEs) are extracted using a frame-semantic parser (FSP).
 Then, FEs are mapped to a table which can then be queried to gather evidence for the claim.
@@
 3 Fact-Checking Paradigm
-We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment.
-Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.” color=yellow]Make it clear in the figure that “Occupy Rank” is the frame.
-Also mention it in the figure caption.
-color=yellow]Why is “life-expectancy” schema listed twice, with slight difference (w/ and w/o measure)?
-color=yellow]In figure 1, spell out FSP as “Frame Semantic Parsing (FSP)” color=yellow]In Figure 1, the yellow boxes are procedures.
-Other colors are data/concept.
-But they are all represented as boxes.
-In addition to color, can we further differentiate them using something else?
-color=yellow]Figure 1 and this section may appear inconsistent.
-“claim understanding”, “evidence retrieval”, “claim-evidence aligenment” is no where to be found in the figure.
-On the other hand, isn’t not clear table mapping, table querying, FSP, fact verfication belong to which of the 3 steps.
-It is like we have presented two alternative naming systems or workflows.
+We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and fact verification.
+Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.”
 3.1 Claim Understanding
@@
 Frame-semantic parsing Das et al.
-(2014) enables the extraction of frames and their frame elements from claims.color=green]can make the concept of FSP more clear.
-Even the next sentence (“identifying frames and … FEs …”) kind of defines it and can be used here instead.
-color=green]Need references for FSP By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification.
-For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.color=green]Be consistent.
-The frame name in the figure doesn’t use underscore.
-Frame-semantics not only facilitates claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources.
+(2014) enables the extraction of frames and their frame elements from claims.
+By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks such as evidence retrieval and verification.
+For example, in Figure 1, the Occupy Rank frame is evoked, which characterizes the claim in terms of an Item occupying a certain Rank along a Dimension.
+This semantic structure provides useful context that can support the system’s interpretation and downstream processing of the claim.
+Frame-semantics not only facilitates claim understanding but also guides the fact-checking process by linking frame elements to appropriate data sources.
 While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables (i.e., congress members and bills), other frames may require a predicted mapping based on the text.
-Figure 1 illustrates an example of this predicted mapping.color=yellow]we can find ways to articulate direct vs.
-predicted mapping more clearly.
-Probably won’t have time before deadline.
+Figure 1 illustrates an example of this predicted mapping.
 3.2 Data Collection and Evidence Retrieval
@@
 Evidence retrieval is guided by the frame elements identified in the claim understanding phase.
+We have manually defined the frame elements used for evidence retrieval for the frames studied in our dataset in Table 6 in the Appendix.
 These frame elements act as filtering conditions for querying the appropriate tables or data sources.
 For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics.
-By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy.
-Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by querying large text corpora instead of a database.
-color=yellow]If we don’t have experiments to support this, we don’t want to discuss it here.
-A future work or discussion section at the end is the right place.
-3.3 Claim-Evidence Alignment
-The final step involves aligning the retrieved evidence with the original claim.
-Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context.
-By only using specific spans of the claim, the system can align the claim and evidence more precisely.
-This approach also enhances explainability, as the system can point to the text used for evidence retrieval.
-Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim.
-For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best.
-While this claim is not entirely accurate, its underlying premise holds true—Japan has an exceptionally high life expectancy.
-As a result, the model would predict it as mostly true.
-color=yellow]This is an intriguing example.
-But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it?
-— this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
+By focusing on specific spans of the claim (i.e., the Dimension frame element “life expectancy”), this method minimizes confusion and enhances retrieval accuracy.
+3.3 Fact Verification
+The final step in our pipeline involves assessing the consistency between the retrieved evidence and the original claim.
+Verifying the truthfulness of a claim requires not only synthesizing the retrieved evidence but also understanding how it relates to the claim and its underlying intent.
+For instance, in Figure 1, the claim states that Japan has the second-best life expectancy, while the data shows it actually has the highest.
+Although the claim is not factually precise, its core premise—that Japan has an exceptionally high life expectancy—remains valid.
+A robust fact verification model should ideally classify such a claim as true or mostly true.
+Modern LLMs are well-suited for capturing such nuances, which is why we adopt an LLM-based approach for this step.
 While fine-tuned approaches have shown slightly better performance on table-based fact-checking Ye et al.
-(2023), we employ an LLM-based approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions.
-Because this step is not a key focus of our work, we leave fact verification model-agnostic to allow future improvements.
-Thus, we can replace the current LLM with any fact verification model which takes structured evidence and a claim to perform fact verification.
+(2023), we employ an LLM-based fact verification approach due to their ability to handle novel inputs without needing to retrain and for their ability to provide natural language explanations of predictions.
+Importantly, since fact verification is not the primary focus of our work, we keep this component model-agnostic.
+That is, our system can accommodate any fact verification model that accepts structured evidence and a claim as input.
+This design choice allows for easy substitution with future state-of-the-art models as they become available.
 4 Fact-Checking Case Studies
 In this section we provide two case studies which utilize our proposed paradigm.
-color=yellow]This is really an important ToDo item. Untify the two case studies in the same terminology framework.
-color=yellow]Another important ToDo item to strengthen the paper: incorporate more case studies.
 4.1 Voting Records
@@
 congressional voting records from the Congress GitHub Repository.
-[^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls.
-Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.color=yellow]We are mentioning 93rd, 101st, and 1st.
-If we only have votes since 101st Congress, why did we include 93rd Congress or even the 1st Congress?
-At least mentioning statistics for them becomes less relevant.
-— this is the data we have, it wouldn’t really make sense to say we only have data from the 101st since we technically have members from before
+[^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd to the 117th Congress.
+It also contains biographical information for 12,677 unique members of Congress dating back to the 1st Congress.
+However, comprehensive roll call voting records are only available from the 101st to the 117th Congress, totaling 7,195,798 individual votes across 22,447 roll calls.
+Each member of Congress cast an average of 4,230 votes during this period.
+Accordingly, our analysis is restricted to data from the 101st Congress onward, which reflects the full span of available voting records.
 PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims.
@@
 To verify voting records, mapping the Agent FE to the correct member of Congress is essential.
-We use SQL queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4).
+We use database queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4).
 In cases of name ambiguity we default to the more recent member.
@@
 To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members.
-Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz.
-We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives when necessary.
+Similarly, some members go by shortened or preferred names, such as “Joe Biden” instead of “Joseph Biden” or “Ted Cruz” instead of “Rafael Edward Cruz”.
+We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives.
 Bill matching.
@@
 Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval.
-To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
+To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the style of language used in the claim and the bill differs significantly.
 Verifying alignment of claim to bill.
 Determining whether a claim is refuted or supported by evidence presents several challenges.
-First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (for or against).
+First, the system cannot rely solely on the vote (Yea or Nay) and the Position FE (for or against).
 [^6] Note that Position FE is extracted from claims by our frame semantic parsing model.
@@
 Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications.
-color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”.
-Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
+To address these concerns, we instruct a large language model (LLM) with these challenges in mind using the prompt in Appendix E.2.
 Fact verification.
@@
 The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics which serve as a strong focal point to explore a wider range of semantic frames.
+For example, there are statistics related to health (including healthcare coverage, health risk factors, pharmaceutical markets, etc.), financial information (including pension assets, stocks and investments, employment earnings, unemployment rates, etc.), scientific data (air pollution, greenhouse gas emissions, exposure to droughts/floods/wildfires, energy consumption, plastic leakage, etc.), and many more topics.
+With this dataset and the wider range of frames, we drastically expand both the scope of topics (leveraging the OECD data) and the types of claims (leveraging the frames) our system can check.
 For the OECD case study, we collected all of the data tables available on the OECD Data Explorer.
 [^7] https://data-explorer.oecd.org/ We constructed an SQL database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totaling over 4.1 billion cells.
+To evaluate our system, like the previous case study, we constructed an evaluation dataset by extracting PolitiFact fact-checks before April 2022 that cite oecd.org in their fact-check sources.
+After manual verification, we collected 68 fact-checks, along with their corresponding OECD data table, to form the evaluation set.
 Relevant table identification.
@@
 To represent the table, we encode the table’s name and description as text.
-To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element from the claim.
+To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element (shown for each frame in Table 6) from the claim.
 Querying retrieved tables.
@@
 To do this, we encode each table’s name, all of its columns, and several representative sample values from each column into text.
-We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table accordingly.
-We choose this approach, as opposed to standard text-to-SQL approaches because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
+We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table(s) accordingly.
+We choose this approach, as opposed to standard text-to-SQL approaches Lei et al.
+(2024) because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
 We represent the table schema as SQL code, following best practices from previous work Gao et al.
@@
 Our implementation is similar to Nahid and Rafiei (2024); however, we address critical limitations to support the high-volume OECD data.
-Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values.
-For other columns we use the same RoBERTa-based semantic similarity model to select the top ten most relevant values based on the claim.
-These queries often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim.
+Encoded columns with fewer than five distinct values include all values, while columns with more than 100 distinct values are randomly sampled for ten representative values.
+For other columns we use the same RoBERTa-based semantic similarity model from the relevant table identification to select the top ten most relevant values based on the claim.
+Querying the databases often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim.
 The prompts for this process can be found in Appendix E.3.
@@
 Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in their dataset along with the newly-defined frames.
-These frames can be found in Table 6 in the Appendix.
+These frames, along with the frame elements used for identifying relevant tables, can be found in Table 6 in the Appendix.
 PolitiFact fact-checks.
 For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles (i.e., fact-checks) collected as of April 2022.
-Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources.
+Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources used in the fact-check.
 To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we manually identified 79 fact-checks that cite congress.gov in the sources and evoke the Vote frame.
@@
 OECD FSP OECD 0.742 0.873
-Table 1: Performance of frame-semantic parsing model on fact-checking frames alongside simple baseline.
+Table 1: Performance of frame-semantic parsing model on fact-checking frames compared with LLM and random choice baselines.
 Frame Samples (%)
@@
 None 12 (0.1%)
-Table 2: Distribution of semantic frames in PolitiFact.
+Table 2: Distribution of semantic frame predictions in fact-checking articles using GPT-4o-mini.
 5.2 Frame-Semantic Parsing
@@
 The prompt used for this model can be found in Appendix E.1.
-We evaluated each model using exact match accuracy on frame identification and argument identification.
+We evaluated each model using exact match accuracy on frame identification and argument identification (predicting the frame elements of the evoked frame).
 These results are presented in Table 1.
@@
 To assess the broader applicability of our frame-semantics-based fact-checking system, we conducted a comprehensive coverage analysis.
+For this analysis, we used the frames discussed in Arslan et al.
+(2020), including the 11 newly defined frames from their work.
 Because the annotations available for the fact-checking frames are limited, we employed a zero-shot GPT-4o-mini model (detailed in Appendix E.1) to identify frames evoked in the PolitiFact fact-check corpus.
 Claims that did not trigger any of our studied frames were classified as “None.” The distribution of frames is presented in Table 2.
-Our findings show that GPT-4o-mini may predict more frames than there actually are.
-This is because the annotated dataset evokes an average of 1.1 frames per claim while GPT-4o-mini predicted an average of 2.1 frames per claim.
-We believe this is because GPT-4o-mini predicts frames for lexical units–words which evoke frames–that are not defined by Arslan et al.
-(2020).color=yellow]This last sentence appears to suggest GPT can find frames there were missed in Fatma’s dataset.
-But why did you say it “predict more than there actually are”?
-— This isn’t really known 100% the cause could be that the model over-predicts in a bad way or it could be because the model recognizes frames for more LUs than are defined.
-It seems to be the latter based on what ive seen.
+Applying GPT-4o-mini resulted in 45,772 predicted frames, indicating that it may over-predict relative to the annotated data.
+While the gold dataset averages 1.1 frames per claim, GPT-4o-mini predicts 2.1 frames on average.
+This likely stems from GPT-4o-mini assigning frames to lexical units—words that can evoke frames—that were not defined by Arslan et al.
+(2020).
 One indication of this can be seen with the Vote frame.
@@
 BM25 Vote 0.048
+distilroberta-base Vote 0.114
+all-mpnet-base-v2 Vote 0.115
 roberta-base-v2 Vote 0.131
+multi-qa-mpnet-base Vote 0.143
+stella_en_400M_v5 Vote 0.144
 stella_en_1.5B_v5 Vote 0.144
 distilbert-tas-b Vote 0.165
+stella_en_1.5B_v5 OECD 0.126
 BM25 OECD 0.323
@@
 text-embedding-3-large OECD 0.610
+roberta-base-v3 OECD 0.642
 roberta-base-v2 OECD 0.726
@@
 To retrieve relevant information for fact-checking, we first evaluated different semantic similarity models.
-Table 3 presents the embedding models we compared, including Stella Zhang et al.
+We experimented with several models from the MTEB leaderboard.
+[^9] https://huggingface.co/spaces/mteb/leaderboard Table 3 presents the best-performing models, including Stella Zhang et al.
 (2025), RoBERTa Liu et al.
 (2019), DistilBERT-TAS-B Hofstätter et al.
-(2021), OpenAI’s text-embedding-3, 999 https://platform.openai.com/docs/models and BM25.
-101010 https://github.com/dorianbrown/rank_bm25 color=yellow]In Table 4, font size of different models are different; some R@K is italic, others not.
-Table 1 and Table 3/4 use different font styles.
+(2021), OpenAI’s text-embedding-3, 101010 https://platform.openai.com/docs/models and BM25.
+111111 https://github.com/dorianbrown/rank_bm25
 Our analysis revealed that different models excel in specific verification contexts.
@@
 We explored two different methods for constructing retrieval queries: using the complete claim text and using only the frame elements (FEs) extracted based on the evoked frame (detailed in Appendix B).
+For each frame, we selected specific frame elements (Table 6) which are most relevant for evidence retrieval.
 We evaluated these query representations using the best-performing similarity models identified in Table 3.
@@
 We evaluated our fact verification system by comparing its predicted verdicts against the ground truth verdicts from PolitiFact fact-checks (Table 5).
-For consistency in our analysis, we consolidated PolitiFact’s "pants on fire" rating with "false".
+For consistency in our analysis, we consolidated PolitiFact’s “pants on fire” rating with “false”.
 To assess whether our system genuinely relies on retrieved evidence rather than pre-existing knowledge, we established a baseline using GPT-4o without access to any external data (GPT-4o Naive).
@@
 Another limitation is that our system’s effectiveness is constrained by the availability of reliable, manually curated data.
-For claims that lack relevant data in our databases (e.g., OECD tables or U.S.
+For claims that lack relevant data in our databases (i.e., OECD tables or U.S.
 congressional records), the system cannot retrieve evidence, which limits its fact-checking capabilities.
 This limitation demonstrates the importance of expanding data sources and improving coverage in future iterations of the system.
+Finally, the method relies on existing FrameNet frames.
+FrameNet versions for languages other than English are often incomplete or entirely unavailable.
+As a result, the current method may not be effective for fact-checking in non-English languages.
 Ethics and Risks
@@
 Similarly, the system could be exploited by adversaries who craft ambiguous or misleading claims designed to confuse AI, leading to manipulated fact-checks.
-Lastly, the over-dependence on specific sources like PolitiFact or congress.gov could limit the tool’s scope.
+Lastly, the over-dependence on specific sources such as PolitiFact or congress.gov could limit the tool’s scope.
 Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
+Acknowledgments
+This work is partially supported by the National Science Foundation under Grant 2346261.
+We also extend our gratitude to the Texas Advanced Computing Center (TACC) for providing compute resources used in this work’s experimentation.
 Appendix A Data Quality
@@
 No disagreements occurred during annotation.
-Model Parameters URL
-msmarco-distilbert-base-tas-b 66M msmarco-distilbert-base-tas-b
-msmarco-roberta-base-v2 125M msmarco-roberta-base-v2
-msmarco-roberta-base-v3 125M msmarco-roberta-base-v3
-stella_en_1.5B_v5 1.5B stella_en_1.5B_v5
-text-embedding-3-large unknown text-embedding-3-large
-Table 7: Models and their corresponding Hugging Face URLs.
+Model Parameters
+msmarco-distilbert-base-tas-b 66M
+msmarco-roberta-base-v2 125M
+msmarco-roberta-base-v3 125M
+stella_en_1.5B_v5 1.5B
+text-embedding-3-large unknown
+Table 7: Models and their parameter sizes.
 Voting claims.
@@
 The OECD database requires roughly 85GB of storage and the voting records requires 1GB of storage.
-color=cyan]we don’t train any other models.
 C.2 Software and Dependencies
@@
 Frame-semantic parsing.
-Our frame-semantic parsing model is an adaptation of AGED, augmented with the frame classification layer and strategy from Devasier et al.
-(2024).
-We fine-tuned this model using a joint loss function to improve frame prediction accuracy.
-Instead of BERT, as originally used in AGED, we employ RoBERTa for its improved contextual embeddings, which have been shown to enhance performance on frame-based semantic parsing tasks.
+Our frame-semantic parsing model builds on AGED, which achieves state-of-the-art performance in identifying the spans of frame elements for a given frame in a sentence.
+While AGED focuses on span detection, Devasier et al.
+(2024) propose a complementary training strategy that improves frame identification by incorporating additional negative examples during training.
+Both models use similar underlying architectures, which allowed us to integrate their strengths.
+Specifically, we augmented AGED with an additional classification layer to predict the correct frame and adopted the robust training procedure introduced by Devasier et al.
+For our base encoder, we used RoBERTa instead of BERT, as it provides stronger contextual representations.
+The resulting model demonstrated strong empirical performance, so we retained this architecture for all downstream tasks.
 Voting agent matching.
-Our agent matching function retrieves the BioGuide ID of a political figure from a database.
-It extracts the agent’s name using frame element spans (FEs) and checks if it is a known presidential nickname, replacing it with the full name if applicable.
-If not, it uses a SpaCy NER model to extract PERSON named entities.
-The function first attempts an exact match in the members table, then a fuzzy match using SQL’s LIKE operator.
+Our agent matching function retrieves the BioGuide ID  [^13] https://www.congress.gov/help/field-values/member-bioguide-ids of a political figure from the voting records database.
+It extracts the agent’s name using frame element spans and checks if it is a known political nickname, replacing it with the full name if applicable.
+If not, it uses a SpaCy NER model [^14] https://spacy.io/api/entityrecognizer to extract PERSON named entities.
+The function first attempts an exact match then a fuzzy match using SQL’s LIKE operator if the exact match fails.
 If neither succeeds, it constructs a multi-condition query, incorporating nickname expansions where applicable.
@@
 Table 8 presents two sample claims from the voting dataset along with their top three retrieved bills.
-For each retrieved bill, we include the computed vote-claim alignment score, which indicates the relevance of the bill to the claim.
-This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.color=red]I don’t any score in the table.
+For each retrieved bill, we include the computed vote-claim alignment classification (one of Supports, Refutes, Irrelevant).
+This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.
 D.2 Fact Verification Predictions
@@
 For voting-related claims, we provide two examples where our model predicts the veracity of claims based on retrieved bills.
-Table 8 includes the claims, retrieved evidence, and the final prediction score.color=red]no score
+Table 8 includes the claims, retrieved evidence, and the final prediction class.
 OECD claims.